### v1

# Structural properties of Krylov subspaces and applications to unbounded self-adjoint operators

Abstract.
This paper presents a study of the inherent structural properties of Krylov subspaces, in particular for the self-adjoint class of operators, and how they relate with the important phenomenon of ‘Krylov solvability’ of linear inverse problems. Owing to the complexity of the problem in the unbounded setting, recently developed perturbative techniques are used that exploit the use of the weak topology on $\mathcal{H}$. We also make a strong connection between the approximative properties of the Krylov subspace and the famous Hamburger problem of moments, in particular the determinacy condition.
1. Introduction
‘Krylov solvability’ of an inverse linear problem is an operator-theoretic phenomenon, with deep implications for applied and theoretical numerical analysis, that has garnered recent attention [7, 8, 4, 5, 6, 10]. An inverse linear problem on a Hilbert space $\mathcal{H}$ is formulated as
(1.1) $Af=g\,,\quad g\in\mathrm{ran}A\,,$
where $A:\mathcal{H}\to\mathcal{H}$ is a closed and densely defined operator, and $f$ is a solution to the above problem. As $g\in\mathrm{ran}A$ we call (1.1) ‘solvable’. If, additionally, $A$ is injective, we call (1.1) ‘well-defined’; and if furthermore we suppose that $A^{-1}$ is continuous and everywhere defined on $\mathcal{H}$, the problem (1.1) is called ‘well-posed’.
In practical circumstances arising from numerical analysis, one searches for a solution(s) $f$ to (1.1) within the celebrated Krylov subspace by using any member of the large family of Krylov algorithms. Therefore, it is critical to know under what circumstances does a solution belong (or fail to belong) to the closure of the Krylov subspace. Such an understanding is important before the actual algorithm is selected and run on the particular problem in order to decide whether the treatment by a particular algorithm has the possibility to yield a correct approximation to a solution of (1.1).
More precisely, the Krylov subspace is constructed using the operator $A$ and known datum vector $g$ in the following manner
(1.2) $\mathcal{K}\left(A,g\right):=\mathrm{span}\{A^{k}g\,|\,k\in\mathbb{N}_{0}\}\,,$
or in other words, $\mathcal{K}\left(A,g\right)$ is the space of all the polynomials in the operator $A$ applied to the vector $g$. Definition (1.2) only makes sense when $g$ is an $A$-smooth vector, i.e., $g\in C^{\infty}(A):=\cap_{n\in\mathbb{N}_{0}}\mathcal{D}(A^{n})$. In purely operator-theoretic jargon, $\mathcal{K}\left(A,g\right)$ is known as the cyclic subspace of the operator $A$ with respect to the vector $g$.
If the inverse linear problem exhibits the existence of a solution in the closure of $\mathcal{K}\left(A,g\right)$, we call such a solution a Krylov solution, and we say that (1.1) is Krylov solvable. Until recently [8, 4, 5, 6], limited attention has been given to this topic, and mainly in the context of studying the convergence of specific algorithms (e.g., GMRES, CG, LSQR, etc.) to a solution of a inverse linear problem typically with specific assumptions on the operator (e.g., $A$ is bounded, positive, purely discrete spectrum, etc)–see [19, 20, 12, 14, 15, 3, 2, 9].
The recent studies [8, 4] have indicated a strong link between the Krylov solvability properties of an inverse linear problem and the inherent structural properties of the Krylov subspace, even though the latter may be difficult to access and therefore analyse. In particular, a construct known as the ‘Krylov intersection’ has been identified as a critical structure to determine whether or not the inverse linear problem is Krylov solvable (see [8, 4]). In the bounded operator setting, the analysis of structural properties of Krylov subspaces simplifies owing to the absense of domain issues: indeed any vector $x\in\mathcal{H}$ is in the domain of the operator, and therefore we do not have to account for a possible domain issues when studying closed subspaces under the action of the operator.
Though the structural properties of Krylov subspaces in the unbounded setting and their link to the Krylov solvability of the linear inverse problem has already been discussed in a previous study [4], here we take an analogous route by changing the topology of the closure of the Krylov subspace by using the graph norm. This permits us to avoid complications that arise in establishing links between Krylov solvability and the structural properties of the Krylov subspace as we focus on the class of closed operators acting on a Hilbert space $\mathcal{H}$. Such complications naturally arise when using the ambient topology of $\mathcal{H}$ due to domain issues when we consider closures of subsets of the domain of an unbounded operator. More specifically, the complications in [4] manifest in the additional assumption that certain solution(s) to (1.1), when projected onto $\overline{\mathcal{K}\left(A,g\right)}$, should remain within the domain of the operator. In both practical and analytical situations, this projection criterion may be difficult to verify.
In this article for the first time we unmask also the strong link between the structural properties of Krylov subspaces arising from self-adjoint operators and the famous problem of moments on the real line, or Hamburger moment problem, and orthogonal polynomials. The Hamburger moment problem and orthogonal polynomials is a classical area of analysis that also has intimate links to the theory of symmetric operators and their self-adjoint extensions (see [22] for an excellent overview). This link between Krylov subspaces generated by self-adjoint operators and the Hamburger moment problem turns out to be of paramount importance for describing the approximation properties of the Krylov subspace, in particular when it is isometrically isomorphic to an $L^{2}$ measure space. Indeed, the existence of a unique solution to the Hamburger moment problem indicates that the Krylov subspace possesses good approximation properties, as outlined in Remark 4.10. These approximation properties permit us to investigate further the structural properties of Krylov subspaces, and in particular the notion of Krylov intersection as well as the so-called ‘Krylov core condition’ (see [4] for details).
This paper is split into several parts. Section 2 deals with the reformulation of the Krylov intersection and the Krylov solvability of the inverse linear problem under a change in topology to that of the graph norm of the operator on its domain. Sections 3.1 and 3.2 review structural notions of Krylov subspaces and the concept of the weak gap, respectively. Section 3.3 reviews the convergence properties of linear subspaces under the weak gap, as well as presenting several new results. Section 4 is split into two main parts: Section 4.1 reviews the basics of the Hamburger moment problem, and Section 4.2 uses this theory to uncover structural properties of Krylov subspaces arising from self-adjoint operators. Lastly, in Section 5 we close the study with a perturbative analysis of Krylov subspaces as well as an analysis of the Krylov solvability arising from self-adjoint inverse linear problems.
1.1. Notation
Throughout this note, the operator $A:\mathcal{H}\to\mathcal{H}$ is a closed operator on a Hilbert space $\mathcal{H}$ with domain $\mathcal{D}\left(A\right)$. As $A$ is closed, the space $V:=(\mathcal{D}\left(A\right),\left\|\cdot\right\|_{A})$ is complete, where
(1.3) $\left\|x\right\|_{A}^{2}:=\left\|x\right\|_{\mathcal{H}}^{2}+\left\|Ax\right\|_{\mathcal{H}}^{2}\,,$
[…]
(1.4) $\left\langle x,y\right\rangle_{V}:=\left\langle x,y\right\rangle+\left\langle Ax,Ay\right\rangle$
for $x,y\in\mathcal{D}\left(A\right)$, we define the appropriate scalar product on the space $V$ that induces $\left\|\cdot\right\|_{A}$, thereby making $V$ a Hilbert space.
Throughout the note, given any $M\subset\mathcal{H}$, we denote by $\perp$ the standard orthogonal complement in $\mathcal{H}$,
$M^{\perp}:=\{x\in\mathcal{H}\,|\,\left\langle x,y\right\rangle=0\,,\,\forall\,y\in M\}\,,$
and under the further condition that $M\subset V$, by $\perp_{V}$ we denote the space
$M^{\perp_{V}}:=\{x\in V\,|\,\left\langle x,y\right\rangle_{V}=0\,,\,\forall\,y\in M\}\,.$
We also denote $\overline{\,\cdot\,}$ to denote closure in the ambient Hilbert space $\mathcal{H}$ with the norm $\left\|\cdot\right\|_{\mathcal{H}}$, and ¯$\overline{\,\cdot\,}^{V}$ to denote the closure in $V$ with the norm $\left\|\cdot\right\|_{A}$. We use $\left\|\cdot\right\|_{\mathrm{op}}$ to denote the standard operator norm from $\mathcal{H}$ to $\mathcal{H}$.
It is well known that the operator $A$, when viewed as acting from the space $V$ (a Hilbert space with stronger topology than its embedding space $\mathcal{H}$) to $\mathcal{H}$, is a bounded operator, i.e., $A\in\mathcal{B}(V,\mathcal{H})$.
2. The Krylov Intersection
In this section we develop the Krylov intersection using the space $V$. We therefore avoid the necessity that projection onto the closed Krylov subspace (in $\mathcal{H}$) of at least one solution to the inverse linear problem remain in the domain of the operator in order to establish Krylov solvability of the inverse linear problem. To clarify the precise meaning of this point, we state the relevant proposition from [4].
Proposition 2.1 (Proposition 6.5, [4]).
For a given Hilbert space $\mathcal{H}$ let $A$ be a densely defined and closed operator on $\mathcal{H}$, let $g\in\mathrm{ran}A\cap C^{\infty}(A)$, and let $f\in\mathcal{D}(A)$ saitsfy $Af=g$. Assume furthermore that
  • (a)
    $A(\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A))\subset\overline{\mathcal{K}\left(A,g\right)}$,
  • (b)
    $P_{\mathcal{K}}f\in\mathcal{D}(A)$, where $P_{\mathcal{K}}$ is the orthogonal projection operator onto the subspace $\overline{\mathcal{K}\left(A,g\right)}$, and
  • (c)
    $\overline{\mathcal{K}\left(A,g\right)}\cap A(\mathcal{K}\left(A,g\right)^{\perp}\cap\mathcal{D}(A))=\{0\}$.
Then there exists $f_{\circ}\in\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A)$ such that $Af_{\circ}=g$.
Proposition 2.1 gives conditions that guarantee the Krylov solvability of the inverse linear problem in the unbounded setting, yet there is the rather strong requirement (b) that the projection onto $\overline{\mathcal{K}\left(A,g\right)}$ of at least one solution to the inverse linear problem remain within the domain of the operator $A$; something that in practice is difficult to verify. The advantage of the theory in this present article is that this assumption (b) in Proposition 2.1 is bypassed altogether.
We now proceed to prove some technical propositions and a lemma that will be needed in proving one of the main results of this section, namely Theorem 2.6.
Proposition 2.2.
Let $A:\mathcal{H}\to\mathcal{H}$ be a closed linear operator, and $g\in\mathcal{C}^{\infty}\left(A\right)$. Then
(2.1) $A\overline{\mathcal{K}\left(A,g\right)}^{V}\subset\overline{\mathcal{K}\left(A,g\right)}\,.$
Proof.
Indeed,
$A\mathcal{K}\left(A,g\right)\subset A\overline{\mathcal{K}\left(A,g\right)}^{V}\subset\overline{A\mathcal{K}\left(A,g\right)}\subset\overline{\mathcal{K}\left(A,g\right)}\,,$
the second inclusion being a consequence of the continuity of A as an operator from $V$ to $\mathcal{H}$ (see [18, Theorem 18.1]). ∎
Lemma 2.3.
Let $A:\mathcal{H}\to\mathcal{H}$ be a closed linear operator, and let $0\in\rho(A)$. Then, in addition to $A^{-1}\in\mathscr{B}(\mathcal{H})$, we have $A^{-1}\in\mathcal{B}(\mathcal{H},V)$. The converse also holds, i.e., if $A^{-1}$ exists and $A^{-1}\in\mathcal{B}(\mathcal{H},V)$, then $0\in\rho(A)$.
Proof.
If $0\in\rho(A)$, we have that $A$ is injective on its domain and surjective. As $A:V\to\mathcal{H}$ is a bounded bijection between $V$ and $\mathcal{H}$, from the open mapping theorem we immediately have that $A$ is a homeomorphism between $V$ and $\mathcal{H}$.
We now show the converse statement. Indeed assuming that $A^{-1}$ exists and is bounded from $\mathcal{H}$ to $V$, we see that
$\left\|A^{-1}\right\|_{\mathrm{op}}=\sup_{\left\|x\right\|_{\mathcal{H}}\leqslant 1}\left\|A^{-1}x\right\|_{\mathcal{H}}\leqslant\sup_{\left\|x\right\|_{\mathcal{H}}\leqslant 1}\left\|A^{-1}x\right\|_{A}=\left\|A^{-1}\right\|_{\mathcal{B}(\mathcal{H},V)}\,.$
Proposition 2.4.
Let $A$ be a closed, injective operator on $\mathcal{H}$, and $g\in\mathrm{ran}A\cap\mathcal{C}^{\infty}\left(A\right)$. Let $f\in\mathcal{D}\left(A\right)$ be the unique solution to $Af=g$.
  • (i)
    If $f\in\overline{\mathcal{K}\left(A,g\right)}^{V}$, then $A\overline{\mathcal{K}\left(A,g\right)}^{V}$ is dense in $\overline{\mathcal{K}\left(A,g\right)}$.
  • (ii)
    Assume further that $0\in\rho(A)$. Then $f\in\overline{\mathcal{K}\left(A,g\right)}^{V}$ if and only if $A\overline{\mathcal{K}\left(A,g\right)}^{V}$ is dense in $\overline{\mathcal{K}\left(A,g\right)}$.
Proof.
We prove (i). Indeed $A\mathcal{K}\left(A,g\right)\subset A\overline{\mathcal{K}\left(A,g\right)}^{V}\subset\overline{\mathcal{K}\left(A,g\right)}$ and $A^{k}g\in A\overline{\mathcal{K}\left(A,g\right)}^{V}$ for all $k\in\mathbb{N}_{0}$ as $f\in\overline{\mathcal{K}\left(A,g\right)}^{V}$. Thus $\mathcal{K}\left(A,g\right)\subset A\overline{\mathcal{K}\left(A,g\right)}^{V}$. Yet, $A\overline{\mathcal{K}\left(A,g\right)}^{V}\subset\overline{\mathcal{K}\left(A,g\right)}$, so the conclusion follows.
To prove (ii) we only consider the ‘if’ implication. Let $A\overline{\mathcal{K}\left(A,g\right)}^{V}$ be dense in $\overline{\mathcal{K}\left(A,g\right)}$, and choose a sequence $(v_{n})_{n\in\mathbb{N}}\subset\overline{\mathcal{K}\left(A,g\right)}^{V}$ such that $\left\|Av_{n}-g\right\|_{\mathcal{H}}\xrightarrow{n\to\infty}0$. As $A^{-1}\in\mathcal{B}(\mathcal{H},V)$ by Lemma 2.3 we see that $(v_{n})_{n\in\mathbb{N}}$ is Cauchy in $V$, and therefore $v_{n}\xrightarrow{n\to\infty}v\in\overline{\mathcal{K}\left(A,g\right)}^{V}$. Furthermore,
$\left\|A^{-1}(Av_{n}-g)\right\|_{A}=\left\|v_{n}-f\right\|_{A}\leqslant\left\|A^{-1}\right\|_{\mathcal{B}(\mathcal{H},V)}\left\|Av_{n}-g\right\|_{\mathcal{H}}\xrightarrow{n\to\infty}0\,.$
Therefore, $v=f$ and so $f\in\overline{\mathcal{K}\left(A,g\right)}^{V}$. ∎
We now define a structure central to the notion of Krylov solvability known as the Krylov intersection. We do not follow the original definition as provided in [4], but instead we propose a new structure in such a way as to avoid certain limitations in [4] that naturally arise as a consequence of the domain of the underlying operator not necessarily defined on all of $\mathcal{H}$.
Definition 2.5.
Let $A$ be a closed linear operator on $\mathcal{H}$, and $g\in\mathcal{C}^{\infty}\left(A\right)$. We define the Krylov intersection $\mathscr{I}_{\mathcal{K}}\left(A,g\right)$ as follows
(2.2) $\mathscr{I}_{\mathcal{K}}\left(A,g\right):=\overline{\mathcal{K}\left(A,g\right)}\cap A\mathcal{K}\left(A,g\right)^{\perp_{V}}\,.$
Theorem 2.6.
Let $A$ be a closed linear operator on $\mathcal{H}$, and let $g\in\mathcal{C}^{\infty}\left(A\right)$ be such that $g\in\mathrm{ran}A$. We have the following.
  • (i)
    If $\mathscr{I}_{\mathcal{K}}\left(A,g\right)=\{0\}$, then there exists a solution $f$ to the inverse linear problem $Af=g$ such that $f\in\overline{\mathcal{K}\left(A,g\right)}^{V}$.
  • (ii)
    If in addition we have that $0\in\rho(A)$, then the solution to the inverse linear problem $Af=g$, $f\in\overline{\mathcal{K}\left(A,g\right)}^{V}$ if and only if $\mathscr{I}_{\mathcal{K}}\left(A,g\right)=\{0\}$.
Remark 2.7.
In comparison to Proposition 2.1 we have avoided the need to ensure that certain solution(s) to the inverse linear problem remain in the domain under the action of the projection operation onto the closure of $\mathcal{K}\left(A,g\right)$ in $\mathcal{H}$. This is overcome by using the graph norm topology (i.e., $V$) and projection operators on $V$ as will become apparent in the course of the following proof.
Proof of Theorem 2.6.
We begin by proving (i). Let $P_{\mathcal{K}}^{V}:V\to V$ denote the orthogonal projection operator (in terms of the scalar product $\left\langle\cdot,\cdot\right\rangle_{V}$) onto the subspace $\overline{\mathcal{K}\left(A,g\right)}^{V}\subset V$. This operator $P_{\mathcal{K}}^{V}$ is continuous from $V$ to $V$. We may always decompose a solution $f$ to $Af=g$ as $f=P_{\mathcal{K}}^{V}f+(\mathbbm{1}_{V}-P_{\mathcal{K}}^{V})f$. Then
$Af=g=AP_{\mathcal{K}}^{V}f+A(\mathbbm{1}_{V}-P_{\mathcal{K}}^{V})f\,.$
As $g\in\overline{\mathcal{K}\left(A,g\right)}^{V}\subset\overline{\mathcal{K}\left(A,g\right)}$ and $AP_{\mathcal{K}}^{V}f\in\overline{\mathcal{K}\left(A,g\right)}$ (Proposition 2.2), we have $A(\mathbbm{1}_{V}-P_{\mathcal{K}}^{V})f\in\overline{\mathcal{K}\left(A,g\right)}$. Yet, we also know that $(\mathbbm{1}_{V}-P_{\mathcal{K}}^{V})f\in\mathcal{K}\left(A,g\right)^{\perp_{V}}$, so that $A(\mathbbm{1}_{V}-P_{\mathcal{K}}^{V})f\in A\mathcal{K}\left(A,g\right)^{\perp_{V}}$ and thus $A(\mathbbm{1}_{V}-P_{\mathcal{K}}^{V})f=0$. This implies $AP_{\mathcal{K}}^{V}f=g$, and therefore $\tilde{f}=P_{\mathcal{K}}^{V}f\in\overline{\mathcal{K}\left(A,g\right)}^{V}$ is a solution to the inverse linear problem.
We now prove the ‘only if’ part of (ii). Let the solution $f$ to $Af=g$ be in $\overline{\mathcal{K}\left(A,g\right)}^{V}$, and suppose that $w\in\mathscr{I}_{\mathcal{K}}\left(A,g\right)$. Then there exists a unique $v\in\mathcal{K}\left(A,g\right)^{\perp_{V}}$ such that $w=Av$. By Proposition 2.4 we know that $A\overline{\mathcal{K}\left(A,g\right)}^{V}$ is dense in $\overline{\mathcal{K}\left(A,g\right)}$, and therefore there exists a sequence $(x_{n})_{n\in\mathbb{N}}\subset\overline{\mathcal{K}\left(A,g\right)}^{V}$ such that $\left\|Ax_{n}-w\right\|_{\mathcal{H}}\to 0$ as $n\to\infty$. Thus
$\left\|x_{n}-v\right\|_{A}=\left\|A^{-1}(Ax_{n}-w)\right\|_{A}\leqslant\left\|A^{-1}\right\|_{\mathcal{B}(\mathcal{H},V)}\left\|Ax_{n}-w\right\|_{\mathcal{H}}\xrightarrow{n\to\infty}0\,.$
Therefore $x_{n}\to v$ in $\left\|\cdot\right\|_{A}$, and this implies $v\in\overline{\mathcal{K}\left(A,g\right)}^{V}$ as well, thus $v=0$. ∎
We shall now develop a numerical indicator to clearly show us whether subspaces are trivially or non-trivially intersecting. We include such a tool as a first step towards a practical numerical indicator of the triviality of $\mathscr{I}_{\mathcal{K}}\left(A,g\right)$ that can be monitored in computations.
Definition 2.8.
Let $M$ and $N$ be subspaces of $\mathcal{H}$. We define the range of ‘separations’ between the subspaces as
(2.3) $\mathcal{S}(M,N):=\left\{\left.\inf_{\begin{subarray}{c}u\in M\,,\\ \left\|u\right\|_{\mathcal{H}}=1\end{subarray}}\left\|u-v\right\|_{\mathcal{H}}\,\right|\,v\in N\,,\,\left\|v\right\|_{\mathcal{H}}=1\right\}\,.$
Remark 2.9.
It is clear that $\mathcal{S}(M,N)\subset[0,\sqrt{2}]$.
Proposition 2.10.
Let $M$ and $N$ be subspaces of $\mathcal{H}$. Then $\overline{M}\cap N=\{0\}$ if and only if $0\notin\mathcal{S}(M,N)$.
Proof.
We begin with the ‘if’ statement first. By contrapositive suppose that there is some $w\in\overline{M}\cap N$ with $\left\|w\right\|_{\mathcal{H}}=1$. Fix $\varepsilon>0$. There exists some $u_{\varepsilon}\in M$ with $\left\|u_{\varepsilon}\right\|_{\mathcal{H}}=1$ such that $\left\|u_{\varepsilon}-w\right\|_{\mathcal{H}}<\varepsilon$ as $w\in\overline{M}$. Therefore, as $\varepsilon$ may be chosen arbitrarily small, it is immediate that
$\inf_{\begin{subarray}{c}u\in M\,,\\ \left\|u\right\|_{\mathcal{H}}=1\end{subarray}}\left\|u-w\right\|_{\mathcal{H}}=0\,,$
which implies that $0\in\mathcal{S}(M,N)$.
We now complete the proof by showing the ‘only if’ part. We also prove this part by contrapositive, so we assume that $0\in\mathcal{S}(M,N)$ and we shall prove that $\overline{M}\cap N\supsetneq\{0\}$. Indeed, take some $v\in N$ such that $\left\|v\right\|_{\mathcal{H}}=1$ and
$\inf_{\begin{subarray}{c}u\in M\,,\\ \left\|u\right\|_{\mathcal{H}}=1\end{subarray}}\left\|u-v\right\|_{\mathcal{H}}=0\,.$
For all $n\in\mathbb{N}$, we choose some $u_{n}\in M$ such that $\left\|u_{n}\right\|_{\mathcal{H}}=1$ and $\left\|u_{n}-v\right\|_{\mathcal{H}}\leqslant\frac{1}{n}$. Then
$\left\|u_{n}-u_{m}\right\|_{\mathcal{H}}\leqslant\left\|u_{n}-v\right\|_{\mathcal{H}}+\left\|u_{m}-v\right\|_{\mathcal{H}}\leqslant\frac{1}{n}+\frac{1}{m}\,,$
meaning that $(u_{n})_{n\in\mathbb{N}}\subset M$ is a Cauchy sequence, and therefore $u_{n}\xrightarrow{n\to\infty}u\in\overline{M}$. Clearly $u=v$, meaning that $v\in\overline{M}\cap N$. ∎
This has an immediate application to the Krylov intersection, as the next theorem shows.
Theorem 2.11.
Let $A:\mathcal{H}\to\mathcal{H}$ be a closed linear operator on Hilbert space $\mathcal{H}$, and $g\in\mathcal{C}^{\infty}\left(A\right)$. Then $\mathscr{I}_{\mathcal{K}}\left(A,g\right)=\{0\}$ if and only if $0\notin\mathcal{S}(\mathcal{K}\left(A,g\right),A\mathcal{K}\left(A,g\right)^{\perp_{V}})$.
Proof.
Use Proposition 2.10, and the fact that $\mathcal{K}\left(A,g\right)$, $A\mathcal{K}\left(A,g\right)^{\perp_{V}}$ are linear subspaces. ∎
Remark 2.12.
$\mathcal{S}(M,N)$ is a wuantity that has the possibility to be monitored at the level of a practical computation. This is a first step to develop such a tool that could be used to monitor the Krylov intersection in a practical way.
3. A review of structural and perturbative properties of Krylov subspaces
In this section, we review some of the important structural and perturbative properties of Krylov subspaces that were revealed in the studies [8, 4, 7] and the monograph [6]. We also present the new Proposition 3.13 linking the convergence of subspaces in the weak-gap metric to the strong operator topology convergence of their orthogonal projection operators.
3.1. Krylov core condition, Krylov reducibility
Here we review structural properties more relevant to this study that are related to Krylov subspaces (aside from the Krylov intersection), as they were introduced in the recent study [4]. The proofs of the various propositions are presented fully in [4].
We begin by defining ‘Krylov reducibility’ in a general unbounded operator setting.
Definition 3.1 (Def.  5.8, [4]).
Let $A$ be a densely defined and closed operator on a Hilbert space $\mathcal{H}$ and let $g\in C^{\infty}(A)$. $A$ is said to be $\mathcal{K}\left(A,g\right)$-reduced in the generalised sense (or informally Krylov reduced) when
(3.1) $\begin{split}A(\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A))\subset\overline{\mathcal{K}\left(A,g\right)}\,,\\ A(\mathcal{K}\left(A,g\right)^{\perp}\cap\mathcal{D}(A))\subset\mathcal{K}\left(A,g\right)^{\perp}\,.\end{split}$
This phenomenon may not always be present. Indeed, already at the level of a bounded everywhere defined operator, the second inclusion does not necessarily hold (see [8] for an example), though the first inclusion is always guaranteed. Also, the first inclusion may fail to happen in the unbounded setting and in such cases we have the notion of ‘Krylov escape’. An explicit example of the Krylov escape phenomenon was constructed in [4, Example 3]. As opposed to the Krylov escape, the orthogonal complement of the Krylov subspace remains invariant under the action of the adjoint in the sense of the following proposition.
Proposition 3.2 (Lemma 5.1, [4]).
Let $A$ be a densely defined and closed operator acting on a Hilbert space $\mathcal{H}$ and let $g\in C^{\infty}(A)$. Then
(3.2) $A^{*}(\mathcal{K}\left(A,g\right)^{\perp}\cap\mathcal{D}(A^{*}))\subset\mathcal{K}\left(A,g\right)^{\perp}\,.$
We now turn to another important structural property called the ‘Krylov core condition’.
Definition 3.3 (Def. 5.4, [4]).
Let $A$ be a densely defined closed operator acting on a Hilbert space $\mathcal{H}$ and let $g\in C^{\infty}(A)$. The operator vector pair $(A,g)$ is said to satisfy the ‘Krylov core condition’ if $\overline{\mathcal{K}\left(A,g\right)}^{V}=\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A)$.
The Krylov core condition turns out to be a sufficient condition to avoid the occurrence of the Krylov escape phenomenon, as the next proposition shows.
Proposition 3.4 (Lemma 5.6 [4]).
Let $A$ be a closed densely defined operator on the Hilbert space $\mathcal{H}$ and $g\in C^{\infty}(A)$. If $(A,g)$ satisfies the Krylov core condition then
$A(\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A))\subset\overline{\mathcal{K}\left(A,g\right)}\,,$
and therefore there is no possibility of ‘Krylov escape’ occurring.
3.2. Weak-gap metric
In this section we review the ‘weak-gap metric’ that has been used to study the effects of perturbations in operator and or datum vector on the Krylov subspace [7].
Naturally, the unit ball $B_{\mathcal{H}}$ of any separable Hilbert space $\mathcal{H}$ is metrisable in the weak topology (see [1, Theorem 3.29] for further details) with a metric $\rho_{w}(x,\,y):=\left\|x-y\right\|_{w}$ such that $\left\|x\right\|_{w}\leqslant\left\|x\right\|_{\mathcal{H}}$. More precisely, for a dense countable collection of vectors $(v_{n})_{n\in\mathbb{N}}\subset B_{\mathcal{H}}$, we have
(3.3) $\left\|x\right\|_{w}:=\sum_{n=1}^{\infty}\frac{1}{2^{n}}|\left\langle v_{n},x\right\rangle_{\mathcal{H}}|\,,$
where $\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$ is the scalar product on $\mathcal{H}$ corresponding to the norm $\|\cdot\|_{\mathcal{H}}$. Therefore the weak convergence of a sequence of vectors $(u_{n})_{n\in\mathbb{N}}\subset B_{\mathcal{H}}$ is tantamount to their convergence in the metric $\rho_{w}$.
The weak-gap metric applies not to individual vectors, but to the class of nonempty weakly closed sets contained in $B_{\mathcal{H}}$
(3.4) $\mathscr{C}_{w}(\mathcal{H}):=\{C\subset B_{\mathcal{H}}\,|\,C\neq\emptyset\,,C\text{ weakly closed}\}\,.$
For $C,D\in\mathscr{C}_{w}(\mathcal{H})$ we have
(3.5) $d_{w}(C,D):=\sup_{u\in C}\inf_{v\in D}\left\|u-v\right\|_{w}\,,$
which then gives us the weak-gap metric on $\mathscr{C}_{w}(\mathcal{H})$
(3.6) $\hat{d}_{w}(C,D):=\max\{d_{w}(C,D),d_{w}(D,C)\}\,.$
We now list some important properties of the weak-gap metric on $\mathscr{C}_{w}(\mathcal{H})$ (without proof).
Theorem 3.5 (Theorem 5.1, [7]).
Let $\mathcal{H}$ be a separable Hilbert space.
  • (i)
    $\hat{d}_{w}$ is a metric on $\mathscr{C}_{w}(\mathcal{H})$.
  • (ii)
    The metric space $(\mathscr{C}_{w}(\mathcal{H}),\hat{d}_{w})$ is complete and compact.
  • (iii)
    If $\hat{d}_{w}(C_{n},C)\xrightarrow{n\to\infty}0$ for some element $C$ and a sequence $(C_{n})_{n\in\mathbb{N}}$ in $\mathscr{C}_{w}(\mathcal{H})$, then
    (3.7) $C=\{u\in B_{\mathcal{H}}\,|\,u_{n}\rightharpoonup u\text{ for a sequence }(u_{n})_{n\in\mathbb{N}}\text{ with }u_{n}\in C_{n}\,\forall\,n\in\mathbb{N}\}.$
  • (iv)
    Suppose $f:\mathcal{H}\to\mathcal{H}$ is a weakly closed and weakly continuous map such that $f(B_{\mathcal{H}})\subset B_{\mathcal{H}}$. If $\hat{d}_{w}(C_{n},C)\xrightarrow{n\to\infty}0$ for an element $C$ and sequence $(C_{n})_{n\in\mathbb{N}}$ in $\mathscr{C}_{w}(\mathcal{H})$, then $\hat{d}_{w}(f(C_{n}),f(C))\xrightarrow{n\to\infty}0$
Before moving on to the application of the weak-gap metric to linear subspaces, we shall list some more properties that are useful later.
Proposition 3.6 (Lemma 5.3, [7]).
Let $C,D,E\in\mathscr{C}_{w}(\mathcal{H})$ for some separable Hilbert space $\mathcal{H}$. Then
(3.8) $d_{w}(C,D)=0\text{ if and only if }C\subset D\,,$
(3.9) $\hat{d}_{w}(C,D)=0\text{ if and only if }C=D\,,$
(3.10) $d_{w}(C,E)\leqslant d_{w}(C,D)+d_{w}(D,E)\,,$
(3.11) $\hat{d}_{w}(C,E)\leqslant\hat{d}_{w}(C,D)+\hat{d}_{w}(D,E)\,.$
3.3. Convergence of linear subspaces in the weak-gap metric
In order to exploit the weak-gap metric for closed linear subspaces of $\mathcal{H}$, we describe convergence on the subsets naturally induced by the corresponding unit balls of the subspaces as elements of $(\mathscr{C}_{w}(\mathcal{H}),\hat{d}_{w})$. More precisely, given closed linear subspaces $M,N$ of a separable Hilbert space $\mathcal{H}$, we identify by definition
(3.12) $\hat{d}_{w}(M,N)\equiv\hat{d}_{w}(B_{M},B_{N})\,,$
with the right side defined by (3.6) and $B_{M}=M\cap B_{\mathcal{H}}$, $B_{N}=N\cap B_{\mathcal{H}}$. We write $\hat{d}_{w}(M_{n},M)\xrightarrow{n\to\infty}0$ to mean that $\hat{d}_{w}(B_{M_{n}},B_{M})\xrightarrow{n\to\infty}0$ for some $M$ and sequence $(M_{n})_{n\in\mathbb{N}}$ closed linear subspaces of $\mathcal{H}$. This naturally provides a metric topology and notion of convergence on the collection of closed subspaces
(3.13) $\mathscr{S}(\mathcal{H}):=\{M\subset\mathcal{H}\,|\,M\text{ closed linear subspace}\}\,,$
thus giving us the metric space $(\mathscr{S}(\mathcal{H}),\hat{d}_{w})$.
We now list the most important properties of $(\mathscr{S}(\mathcal{H}),\hat{d}_{w})$.
Proposition 3.7 (Lemmas 6.1, 6.2, Proposition 6.3, [7]).
Let $\mathcal{H}$ be a separable Hilbert space.
  • (i)
    $(\mathscr{S}(\mathcal{H}),\hat{d}_{w})$ is a metric space.
  • (ii)
    $(\mathscr{S}(\mathcal{H}),\hat{d}_{w})$ is not complete.
  • (iii)
    If $\hat{d}_{w}(M_{n},M)\xrightarrow{n\to\infty}0$ for some $(M_{n})_{n\in\mathbb{N}}$ and $M$ in $\mathscr{S}(\mathcal{H})$, then
    (3.14) $M=\{u\in\mathcal{H}\,|\,u_{n}\rightharpoonup u\text{ for a sequence }(u_{n})_{n\in\mathbb{N}}\text{ with }u_{n}\in M_{n}\text{ for all }n\in\mathbb{N}\}\,.$
Remark 3.8.
Though there is a lack of completeness of $(\mathscr{S}(\mathcal{H}),\hat{d}_{w})$, it turns out that $\hat{d}_{w}$ is still useful as will be revealed in the rest of this article. Moreover, though there are Cauchy sequences $(M_{n})_{n\in\mathbb{N}}$ in $(\mathscr{S}(\mathcal{H}),\hat{d}_{w})$ that do not converge to an element of $\mathscr{S}(\mathcal{H})$, the corresponding unit balls $B_{M_{n}}$ in fact do still converge to some element of $\mathscr{C}_{w}(\mathcal{H})$ in $\hat{d}_{w}$ due to the completeness of $(\mathscr{C}_{w}(\mathcal{H}),\hat{d}_{w})$. In such cases it is clear that the limit in $\mathscr{C}_{w}(\mathcal{H})$ cannot be written as the unit ball of a closed linear subspace of $\mathcal{H}$.
Following this review, we now present some new results of a technical nature related to subspaces and convergence in the weak-gap metric.
Lemma 3.9.
Let $\mathcal{H}$ be a separable Hilbert space, and let $(C_{n})_{n\in\mathbb{N}}$ be a sequence of weakly closed nested sets in $B_{\mathcal{H}}$ such that $C_{n}\subset C_{n+1}$. Then there exists a weakly closed set $C\subset B_{\mathcal{H}}$ such that $C_{n}\xrightarrow{\hat{d}_{w}}C$ as $n\to\infty$.
Proof.
We use the facts that the collection of weakly closed sets in $B_{\mathcal{H}}$ equipped with the metric $\hat{d}_{w}$ is compact (Theorem 3.5(ii)), and both $d_{w}$ and $\hat{d}_{w}$ satisfy the triangle inequality (Proposition 3.6). We also note that for any two weakly closed sets $D,E\subset B_{\mathcal{H}}$, we have $d_{w}(D,E)=0$ if and only if $D\subset E$ (Proposition 3.6).
Let $(C_{n})_{n\in\mathbb{N}}$ be a sequence of weakly closed sets as stated in the lemma. From the fact that $(\mathscr{C}_{w}(\mathcal{H}),\hat{d}_{w})$ is compact, there exists a subsequence $(C_{n_{k}})_{k\in\mathbb{N}}$ that converges to some weakly closed set $C\subset B_{\mathcal{H}}$ in the metric $\hat{d}_{w}$. Let $\varepsilon>0$, so there exists $N_{\varepsilon}\in\mathbb{N}$ such that for all $k,k^{\prime}\geqslant N_{\varepsilon}$
$\hat{d}_{w}(C_{n_{k}},C_{n_{k^{\prime}}})$ $<\varepsilon\,,\text{ and }$
$\hat{d}_{w}(C_{n_{k}},C)$ $<\varepsilon\,.$
Let $N_{\varepsilon}^{\prime}=n_{N_{\varepsilon}}$ and choose any $n\geqslant N_{\varepsilon}^{\prime}$. Then there exist $k,k^{\prime}\geqslant N_{\varepsilon}$ such that $C_{n_{k}}\subset C_{n}\subset C_{n_{k^{\prime}}}$. Using the triangle inequality for $d_{w}$,
$d_{w}(C_{n},C)$ $\leqslant d_{w}(C_{n},C_{n_{k^{\prime}}})+d_{w}(C_{n_{k^{\prime}}},C)$
$=0+d_{w}(C_{n_{k^{\prime}}},C)<\varepsilon\,,$
$d_{w}(C,C_{n})$ $\leqslant d_{w}(C,C_{n_{k}})+d_{w}(C_{n_{k}},C_{n})$
$=d_{w}(C,C_{n_{k}})+0<\varepsilon\,.$
This implies that $\hat{d}_{w}(C,C_{n})<\varepsilon$ for all $n\geqslant N_{\varepsilon}^{\prime}$, and therefore we have the desired convergence. ∎
Lemma 3.10.
Let $\mathcal{H}$ be a separable Hilbert space and let $(C_{n})_{n\in\mathbb{N}}$ be a sequence of weakly closed nested sets in $B_{\mathcal{H}}$ such that $C_{n}\supset C_{n+1}$. Then there exists a weakly closed set $C\subset B_{\mathcal{H}}$ such that $C_{n}\xrightarrow{\hat{d}_{w}}C$ as $n\to\infty$.
Proof.
The proof is identical to Lemma 3.9
Proposition 3.11.
Let $\mathcal{H}$ be a separable Hilbert space. Consider $(M_{n})_{n\in\mathbb{N}}$ a nested increasing sequence of closed linear subspaces such that
(3.15) $M_{n}\xrightarrow{\hat{d}_{w}}M\,,$
for some $M$ a closed linear subspace of $\mathcal{H}$. Then $M_{n}\subset M$ for all $n\in\mathbb{N}$, and $(M_{n}^{\perp})_{n\in\mathbb{N}}$ is a nested decreasing sequence such that
(3.16) $M_{n}^{\perp}\xrightarrow{\hat{d}_{w}}M^{\perp}\,,$
and $M^{\perp}\subset M_{n}^{\perp}$ for all $n\in\mathbb{N}$.
Proof.
Consider $(M_{n})_{n\in\mathbb{N}}$ and $M$ satisfying the hypotheses of the proposition. We know from the convergence of the $M_{n}$’s to $M$ (Theorem 3.5(iii)) that
$B_{M}=\{u\in B_{\mathcal{H}}\,|\,u_{n}\rightharpoonup u\text{ for a sequence }(u_{n})_{n\in\mathbb{N}}\text{ with }u_{n}\in B_{M_{n}}\,\forall\,n\in\mathbb{N}\}\,.$
The nested increasing nature of the $M_{n}$’s means that $M_{n}\subset M_{n+1}$ for all $n\in\mathbb{N}$. Given any $n_{0}\in\mathbb{N}$ and any $u\in B_{M_{n_{0}}}$, for all $n\geqslant n_{0}$ let $u_{n}=u$, and for all $n<n_{0}$ let $u_{n}=0$. Then $(u_{n})_{n\in\mathbb{N}}$ is a sequence such that $u_{n}\in B_{M_{n}}$ for all $n\in\mathbb{N}$ and $u_{n}\rightharpoonup u\in B_{\mathcal{H}}$ as $n\to\infty$. Therefore, $u\in B_{M}$, so $B_{M_{n_{0}}}\subset B_{M}$. As $n_{0}$ is arbitrary, this implies that $B_{M_{n}}\subset B_{M}$ for all $n\in\mathbb{N}$, so that indeed we have the inclusion
$M_{n}\subset M\,,\quad\forall\,n\in\mathbb{N}\,.$
From the nested increasing nature of the $M_{n}$’s it follows that
$M_{n}^{\perp}\supset M_{n+1}^{\perp}\,,\quad\forall\,n\in\mathbb{N}\,,$
and thus $B_{M_{n}^{\perp}}\supset B_{M_{n+1}^{\perp}}$ for all $n\in\mathbb{N}$. By Lemma 3.10 there exists a weakly closed set $C\subset B_{\mathcal{H}}$ such that $\hat{d}_{w}(B_{M_{n}^{\perp}},C)\to 0$ as $n\to\infty$.
We claim that $C=B_{M^{\perp}}$. To this end, let $v\in C$ so that there exists a sequence $(v_{n})_{n\in\mathbb{N}}$ such that $v_{n}\rightharpoonup v$ and $v_{n}\in B_{M_{n}^{\perp}}$ for all $n\in\mathbb{N}$ (Theorem 3.5(iii)). Let $w\in B_{M}$ so that there exists a sequence $(w_{n})_{n\in\mathbb{N}}$ such that $w_{n}\rightharpoonup w$ and $w_{n}\in B_{M_{n}}$ for all $n\in\mathbb{N}$.
We shall construct a subsequence of weakly closed sets $(B_{M_{n_{k}}})_{k\in\mathbb{N}}$ along with a sequence $(w^{\prime}_{n_{k}})_{k\in\mathbb{N}}$ that has the property $w^{\prime}_{n_{k}}\in B_{M_{n_{k}}}$ for all $k\in\mathbb{N}$ and $w^{\prime}_{n_{k}}\to w$ strongly. Indeed, as $w_{n}\rightharpoonup w$, by Mazur’s Theorem [1, Corollary 3.8] there exists a sequence of vectors $(z_{k})_{k\in\mathbb{N}}$ that are made up of convex combinations of the $w_{n}$’s, such that $z_{k}\to w$ strongly. For each $n\in\mathbb{N}$, $\left\|z_{k}\right\|_{\mathcal{H}}\leqslant 1$ as each $z_{k}$ is a convex combination of the $w_{n}$’s and thus remains within the unit ball. As each $z_{k}$ is a convex combination of the $w_{n}$’s, it has the form
$z_{k}=\sum_{n=1}^{N_{k}}t_{n}^{(k)}w_{n}\,,$
where $t_{n}^{(k)}\geqslant 0$, $N_{k}\in\mathbb{N}$, and $\sum_{n=1}^{N_{k}}t_{n}^{(k)}=1$ for all $k\in\mathbb{N}$. Therefore, such a combination exists in the set $B_{M_{N_{k}}}$ given the nested structure of the Krylov subspaces and the convexity of $B_{M_{n}}$ for all $n\in\mathbb{N}$. The indexing of the subsequence $w^{\prime}_{n_{k}}$ is set as follows: let $k=1$ and choose some number $n_{1}\in\mathbb{N}$ such that $z_{1}\in B_{M_{n_{1}}}$. We then set $w^{\prime}_{n_{1}}=z_{1}\in B_{M_{n_{1}}}$. We do the same process with $k=2$, choosing $n_{2}>n_{1}$ which is always possible due to the nested structure of the sequence $(B_{M_{n}})_{n\in\mathbb{N}}$. Continuing this process, we obtain a subsequence of balls $(B_{M_{n_{k}}})_{k\in\mathbb{N}}$ with a sequence $(w^{\prime}_{n_{k}})_{k\in\mathbb{N}}$ such that $w^{\prime}_{n_{k}}\in B_{M_{n_{k}}}$ for each $k\in\mathbb{N}$, and such that $w^{\prime}_{n_{k}}\xrightarrow{k\to\infty}w$ strongly.
Given that $v_{n_{k}}\rightharpoonup v$, and $v_{n_{k}}\in B_{M_{n_{k}}}^{\perp}$ for all $k\in\mathbb{N}$, we see that
$0=\left\langle v_{n_{k}},w^{\prime}_{n_{k}}\right\rangle\xrightarrow{k\to\infty}\left\langle v,w\right\rangle\,,$
implying that $v\perp w$. As the nature of $w$ is arbitrary in $B_{M}$, this proves that $v\in M^{\perp}$, and therefore $C\subset M^{\perp}$.
It is clear from the fact that $M_{n}\subset M$ that we have the inclusion $M_{n}^{\perp}\supset M^{\perp}$ for all $n\in\mathbb{N}$. This implies that $C\supset B_{M^{\perp}}$, and therefore $C=B_{M^{\perp}}$. We have the convergence
$M_{n}^{\perp}\xrightarrow[n\to\infty]{\hat{d}_{w}}M^{\perp}\,,$
thus completing the proof. ∎
Remark 3.12.
The nested structure $M_{n}\subset M_{n+1}$ turned out to be critical for proving the above result, in particular for the use of Mazur’s Theorem.
The next proposition relates convergence of subspaces in the weak-gap metric to the convergence of the orthogonal projections in the strong operator topology.
Proposition 3.13.
Consider $\mathcal{H}$ a separable Hilbert space and let $M$ be a closed linear subspace with $(M_{n})_{n\in\mathbb{N}}$ a sequence of closed linear subspaces that satisfy the following conditions:
  • (i)
    $M_{n}\subset M_{n+1}$ for all $n\in\mathbb{N}$,
  • (ii)
    $M_{n}\xrightarrow{\hat{d}_{w}}M$.
Then $P_{M_{n}}\to P_{M}$ in the strong operator topology, where $P_{M_{n}}$ is the orthogonal projection onto $M_{n}$, and $P_{M}$ is the orthogonal projection onto $M$.
Proof.
We note that under the hypotheses of the proposition $M_{n}^{\perp}\xrightarrow{\hat{d}_{w}}M^{\perp}$ as $n\to\infty$, and also $M_{n}^{\perp}\supset M_{n+1}^{\perp}\supset M^{\perp}$, $M_{n}\subset M$ for all $n\in\mathbb{N}$ by Proposition 3.11.
First we shall establish the claim that for any $x\in\mathcal{H}$, $P_{M_{n}}P_{M_{m}}x=P_{M_{m}}P_{M_{n}}x$ and $P_{M_{n}}P_{M}x=P_{M}P_{M_{n}}x$ for all $n,m\in\mathbb{N}$. Indeed, without loss of generality, we assume that $m>n$, so that $P_{M_{n}}x\in M_{m}$ due to condition (i) and therefore $P_{M_{m}}P_{M_{n}}x=P_{M_{n}}x$. Also, $x=P_{M_{m}}x+(\mathbbm{1}-P_{M_{m}})x$ and $(\mathbbm{1}-P_{M_{m}})x\in M_{m}^{\perp}\subset M_{n}^{\perp}$. Therefore, $P_{M_{n}}x=P_{M_{n}}P_{M_{m}}x+P_{M_{n}}(\mathbbm{1}-P_{M_{m}})x=P_{M_{n}}P_{M_{m}}x$, from which $P_{M_{m}}P_{M_{n}}x=P_{M_{n}}x=P_{M_{n}}P_{M_{m}}x$. Similarly, $P_{M}P_{M_{n}}x=P_{M_{n}}x$ as $P_{M_{n}}x\in M$; and as $(\mathbbm{1}-P_{M})x\in M^{\perp}\subset M_{n}^{\perp}$ we have $P_{M_{n}}x=P_{M_{n}}P_{M}x$. Therefore, $P_{M_{n}}P_{M}x=P_{M}P_{M_{n}}x$ as claimed.
Now we let $x\in B_{\mathcal{H}}$. We may decompose $x$ as
$x=P_{M_{n}}x+P_{M}(\mathbbm{1}-P_{M_{n}})x+(\mathbbm{1}-P_{M})x\,,$
where $P_{M_{n}}x$ is the component in $M_{n}$, $P_{M}(\mathbbm{1}-P_{M_{n}})x$ is the component in $M$ that is orthogonal to $M_{n}$, and $(\mathbbm{1}-P_{M})x$ is the component orthogonal to $M$. Each component is also contained in $B_{\mathcal{H}}$.
From the compactness of $B_{\mathcal{H}}$ in the weak topology [1, Theorem 3.17], there exists a subsequence of $(P_{M}(\mathbbm{1}-P_{M_{n}})x)_{n\in\mathbb{N}}$ that weakly converges to some $v\in B_{\mathcal{H}}$, i.e.,
$P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x\rightharpoonup v\,.$
It is clear that $v\in M$ as the sequence $(P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x)_{k\in\mathbb{N}}\subset M$ and $M$ is weakly closed. Yet, owing to the commutativity of $P_{M}$ and $P_{M_{n_{k}}}$, we also know that $P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x\in M_{n_{k}}^{\perp}$ for all $k\in\mathbb{N}$. Therefore, $v\in M^{\perp}$ from the convergence $M_{n_{k}}^{\perp}\xrightarrow{\hat{d}_{w}}M^{\perp}$ and Proposition 3.7(iii), and thus $v=0$. Furthermore, from the weak convergence $P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x\rightharpoonup 0$, we deduce that
$\left\|P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x\right\|_{\mathcal{H}}^{2}$ $=\left\langle P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x,P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x\right\rangle$
$=\left\langle P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x,(\mathbbm{1}-P_{M_{n_{k}}})x\right\rangle$
$=\left\langle(\mathbbm{1}-P_{M_{n_{k}}})P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x,x\right\rangle$
$=\left\langle P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x,x\right\rangle\xrightarrow{k\to\infty}0\,,$
owing to the commutativity and the fact that $P_{M}$ and $P_{M_{n_{k}}}$ are orthogonal projections. It is then clear that
$P_{M}(\mathbbm{1}-P_{M_{n_{k}}})x\xrightarrow[k\to\infty]{\left\|\cdot\right\|_{\mathcal{H}}}0\,,$
[…]
$P_{M_{n_{k}}}x\xrightarrow[k\to\infty]{\left\|\cdot\right\|_{\mathcal{H}}}P_{M}x\,.$
We now show that the full sequence $(P_{M_{n}})_{n\in\mathbb{N}}$ converges to $P_{M}$ in the strong operator topology. Indeed, let $x\in\mathcal{H}$ and let $\varepsilon>0$. There exists some $N_{\varepsilon,x}\in\mathbb{N}$ such that for all $k,k^{\prime}\geqslant N_{\varepsilon,x}$,
$\left\|P_{M_{n_{k}}}x-P_{M_{n_{k^{\prime}}}}x\right\|_{\mathcal{H}}<\varepsilon\,.$
Choose $n,m\in\mathbb{N}$ such that $n\geqslant n_{N_{\varepsilon,x}+1}$ and without loss of generality, let $m>n$. We choose some $n_{k},n_{k^{\prime}}$ such that $k,k^{\prime}\geqslant N_{\varepsilon,x}$ and $n_{k}<n<m<n_{k^{\prime}}$. Then,
$P_{M_{m}}x-P_{M_{n}}x=(P_{M_{n_{k^{\prime}}}}x-P_{M_{n_{k}}}x)-(P_{M_{n_{k^{\prime}}}}x-P_{M_{m}}x)+(P_{M_{n_{k}}}x-P_{M_{n}}x)\,,$
and
$P_{M_{n_{k^{\prime}}}}x-P_{M_{m}}x=P_{M_{n_{k^{\prime}}}}x-P_{M_{n_{k}}}x+P_{M_{n_{k}}}x-P_{M_{m}}x\,,$
and from the commutativity of the projections and inclusions $M_{m}\subset M_{n_{k^{\prime}}}$, $M_{m}\supset M_{n_{k}}$, we have
$P_{M_{n_{k}}}x-P_{M_{m}}x$ $=P_{M_{n_{k}}}x-P_{M_{n_{k^{\prime}}}}P_{M_{m}}x$
$=P_{M_{n_{k}}}P_{M_{m}}x-P_{M_{n_{k^{\prime}}}}P_{M_{m}}x$
$=(P_{M_{n_{k}}}-P_{M_{n_{k^{\prime}}}})P_{M_{m}}x=P_{M_{m}}(P_{M_{n_{k}}}-P_{M_{n_{k^{\prime}}}})x\,.$
This implies
$\left\|P_{M_{n_{k}}}x-P_{M_{m}}x\right\|_{\mathcal{H}}=\left\|P_{M_{m}}(P_{M_{n_{k}}}-P_{M_{n_{k^{\prime}}}})x\right\|_{\mathcal{H}}\leqslant\left\|(P_{M_{n_{k}}}-P_{M_{n_{k^{\prime}}}})x\right\|_{\mathcal{H}}<\varepsilon\,,$
and therefore
$\left\|P_{M_{n_{k^{\prime}}}}x-P_{M_{m}}x\right\|_{\mathcal{H}}<2\varepsilon\,.$
Using the fact that $P_{M_{n_{k}}}x=P_{M_{n_{k}}}P_{M_{n}}x$ and $P_{M_{n_{k^{\prime}}}}P_{M_{n}}x=P_{M_{n}}x$ owing to the inclusions $M_{n}\supset M_{n_{k}}$ and $M_{n}\subset M_{n_{k^{\prime}}}$ respectively, we see that
$P_{M_{n_{k}}}x-P_{M_{n}}x$ $=P_{M_{n_{k}}}x-P_{M_{n_{k^{\prime}}}}P_{M_{n}}x$
$=P_{M_{n_{k}}}P_{M_{n}}x-P_{M_{n_{k^{\prime}}}}P_{M_{n}}x=P_{M_{n}}(P_{M_{n_{k}}}-P_{M_{n_{k^{\prime}}}}x)\,.$
This implies
$\left\|P_{M_{n_{k}}}x-P_{M_{n}}x\right\|_{\mathcal{H}}\leqslant\left\|P_{M_{n_{k}}}x-P_{M_{n_{k^{\prime}}}}x\right\|_{\mathcal{H}}<\varepsilon\,.$
Putting all these inequalities together
$\left\|P_{M_{n}}x-P_{M_{m}}x\right\|_{\mathcal{H}}<4\varepsilon$
which shows us that $(P_{M_{n}}x)_{n\in\mathbb{N}}$ is a Cauchy sequence. Therefore this sequence has a limit in $\mathcal{H}$ that must be the same as the limit of the convergent subsequence, i.e., $P_{M}x$. Therefore, we have that for all $x\in\mathcal{H}$
$P_{M_{n}}x\xrightarrow[n\to\infty]{\left\|\cdot\right\|_{\mathcal{H}}}P_{M}x\,.$
Remark 3.14.
We note that when subspaces converge in the sense of the Kato gap metric (see [16, Chapter IV, Section 2]) in a Hilbert space, this corresponds to the operator norm convergence of the projections. In comparison, in Proposition 3.13 we manage to achieve the strong operator topology convergence by having convergence in the weak-gap metric.
4. Structural Properties of Self-Adjoint Operators Revisited
We already know from previous analysis [5] that self-adjoint inverse linear problems, under suitable growth conditions on the datum vector $g$, exhibit Krylov-solvability. We aim to show this again in light of the Krylov intersection, and therefore how this interplays with the structural properties of $\mathcal{K}\left(A,g\right)$.
It turns out that the structural properties of Krylov subspaces and spectral properties of the operator $A$ in the self-adjoint setting are very strongly interlinked with the moment problem in one dimension. The moment problem, barring more abstract generalisations, seeks to establish a representation (where one exists) of a linear functional acting on polynomials over $\mathbb{R}$ with integration using a positive Radon measure on $\mathbb{R}$.
This investigation of the link between structural properties of Krylov subspaces and the moment problem in one dimension is a natural follow-up of the study [5]. In [5] we made connections between the convergence of the approximants to a solution to (1.1) from the conjugate gradients algorithm and certain properties of determinacy of the moment problem and properties of orthogonal polynomials on the real line. A very thorough and modern overview of moment problems and orthogonal polynomials is provided in the monograph [22].
A very important structural property of Krylov subspaces generated by a self-adjoint operator is the existence of an isomorphism between $\overline{\mathcal{K}\left(A,g\right)}$ and the measure space $L^{2}(\mathbb{R},\mu_{g}^{(A)})$, where $\mu_{g}^{(A)}(\Omega):=\left\langle E^{(A)}(\Omega)g,g\right\rangle$ and $E^{(A)}$ is the unique projection valued spectral measure for the operator $A$ and $\Omega\subset\mathbb{R}$ a Borel set. The presence of this isomprphism is used in Proposition 4.11 to prove the Krylov-core condition (see [4] for more details).
4.1. The Hamburger Moment Problem
We first discuss important results that stem from the Hamburger moment problem in one dimension which are used heavily in our analysis. The Hamburger moment problem is by now a classical area of mathematics and has been very well studied [22, 11, 17, 24, 23]. We begin by discussing some generalities of the moment problem in an abstract setting before focussing our attention specifically on the Hamburger moment problem.
For the general one dimensional moment problem on $\mathbb{R}$ we consider a linear functional $\mathscr{L}$ on a linear space of continuous functions $V$ over $\mathbb{R}$ and ask: under what conditions does there exist a positive Radon measure $\mu$ on $\mathbb{R}$ such that $\mathscr{L}(f)=\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu(x)$ for all $f\in V$? In the case where there does exist such a Radon measure, we say that $\mathscr{L}$ is a moment functional.
Typically the moment problem is formulated for the case where $V$ is the space of all real valued polynomials over $\mathbb{R}$, and the moment functional $\mathscr{L}$ either: (i) is defined by a sequence of real numbers $s\equiv(s_{n})_{n\in\mathbb{N}_{0}}$ by $\mathscr{L}(x^{n}):=s_{n}$ for all $n\in\mathbb{N}_{0}$, or (ii) it defines a sequence of real numbers $s\equiv(s_{n})_{n\in\mathbb{N}_{0}}$ by $s_{n}:=\mathscr{L}(x^{n})$ for all $n\in\mathbb{N}_{0}$.
Therefore we can restate the one dimensional moment problem as follows: given a sequence of real numbers $s\equiv(s_{n})_{n\in\mathbb{N}_{0}}$ under what conditions does there exist a positive Radon measure $\mu$ on $\mathbb{R}$ such that
(4.1) $s_{n}=\mathscr{L}(x^{n})=\int_{\mathbb{R}}x^{n}\,\mathrm{d}\mu\text{ for all }n\in\mathbb{N}_{0}\,?$
In the case of the existence of such a measure, we say that the sequence $s$ is a Hamburger moment sequence. Moreover we wish to know, in the case of existence of such a measure, whether it is unique–also known as the determinacy of the moment problem. Before delving into this argument we state some necessary definitions and existence results for the Hamburger moment problem.
Definition 4.1.
Let $(s_{n})_{n\in\mathbb{N}_{0}}$ be a sequence of real numbers. We say the the sequence is positive semidefinite if for all $(\xi_{0},\xi_{1},\xi_{2},\dots,\xi_{n})\in\mathbb{R}^{n+1}$ for $n\in\mathbb{N}_{0}$ we have
(4.2) $\sum_{i,j=0}^{n}s_{i+j}\xi_{i}\xi_{j}\geqslant 0\,,$
and we say that the sequence is positive definite if
(4.3) $\sum_{i,j=0}^{n}s_{i+j}\xi_{i}\xi_{j}>0\,,$
for all $(\xi_{0},\xi_{1},\xi_{2},\dots,\xi_{n})\neq 0$.
Theorem 4.2 (Hamburger’s Theorem, Theorem 3.8 [22]).
For a real sequence $s\equiv(s_{n})_{n\in\mathbb{N}_{0}}$ the following are equivalent:
  • (i)
    the sequence $s$ is a Hamburger moment sequence, i.e., there exists a positive Radon measure $\mu$ on $\mathbb{R}$ such that $x^{n}\in L^{1}(\mathbb{R},\mu)$ and
    $s_{n}=\int_{\mathbb{R}}x^{n}\,\mathrm{d}\mu\,,\quad\forall\,n\in\mathbb{N}_{0}\,,$
  • (ii)
    the sequence $s$ is positive semidefinite,
  • (iii)
    the linear functional $\mathscr{L}$ is a positive linear functional on the space of real polynomials, i.e., $\mathscr{L}(p^{2})\geqslant 0$ for all real polynomials $p(x)$.
Remark 4.3.
If we consider any $g\in\mathbb{C}^{\infty}(A)$ for a self-adjoint operator $A$ on Hilbert space $\mathcal{H}$, we see that the positive (and bounded) Radon measure $\mu_{g}^{(A)}(\Omega):=\left\langle E^{(A)}(\Omega)g,g\right\rangle$ on $\mathbb{R}$ generates a linear functional $\mathscr{L}$ on the space of all polynomials over $\mathbb{R}$, $\mathscr{L}(p):=\int_{\mathbb{R}}p(\lambda)\,\mathrm{d}\mu_{g}^{(A)}(\lambda)$ for some polynomial $p$.
Therefore on the subspace of all real polynomials, the functional $\mathscr{L}$ is a positive linear functional, i.e., for any real polynomial $p$ we have $\mathscr{L}(p^{2})\geqslant 0$. By Theorem 4.2 the sequence of real numbers given by $s_{n}:=\mathscr{L}(\lambda^{n})=\int_{\mathbb{R}}\lambda^{n}\,\mathrm{d}\mu_{g}^{(A)}(\lambda)$ is positive semidefinite.
We now move on to the theorem of determinacy of the Hamburger moment problem. Without loss of generality, we suppose that there exists a positive definite sequence of real numbers $s\equiv(s_{n})_{n\in\mathbb{N}}$. From Theorem 4.2 there exists a positive Radon measure $\mu$ on $\mathbb{R}$ such that
$s_{n}=\int_{\mathbb{R}}x^{n}\,\mathrm{d}\mu\,,\quad\forall\,n\in\mathbb{N}_{0}\,,$
and we define the linear functional $\mathscr{L}$ over the space of all polynomials on $\mathbb{R}$ as follows $\mathscr{L}(x^{n}):=\int_{\mathbb{R}}x^{n}\,\mathrm{d}\mu=s_{n}$, for all $n\in\mathbb{N}_{0}$. Then for the space of all polynomials on $\mathbb{R}$ we have that $\mathscr{L}$ defines a scalar product, antilinear in the first entry, in the following way.
(4.4) $\left\langle p,q\right\rangle_{s}:=\mathscr{L}(\bar{p}q)\,,$
for any $p,q$ polynomials on $\mathbb{R}$. We see that given any non-zero polynomial $p$ on $\mathbb{R}$ of degree $n\in\mathbb{N}_{0}$
$\left\langle p,p\right\rangle_{s}=\mathscr{L}(|p|^{2})=\int_{\mathbb{R}}\overline{p(x)}p(x)\,\mathrm{d}\mu=\sum_{i,j=0}^{n}s_{i+j}\bar{c_{i}}c_{j}\,,$
where $p(x)=\sum_{i=0}^{n}c_{i}x^{i}$ and $c_{i}\in\mathbb{C}$ for all $i=1,2,\dots,n$ and $(c_{1},c_{2},\dots,c_{n})\neq 0$. Thus, owing to (4.3) we see that $\left\langle p,p\right\rangle_{s}>0$. Indeed,
$\Re\sum_{i,j=0}^{n}s_{i+j}\bar{c_{i}}c_{j}=\sum_{i,j=0}^{n}s_{i+j}\bar{c_{i}}c_{j}\,,$
and
$\Re(\bar{c_{i}}c_{j})=\Re(c_{i})\Re(c_{j})+\Im(c_{i})\Im(c_{j})\,,$
so that
$\sum_{i,j=0}^{n}s_{i+j}\bar{c_{i}}c_{j}=\sum_{i,j=0}^{n}s_{i+j}\Re(c_{i})\Re(c_{j})+\sum_{i,j=0}^{n}s_{i+j}\Im(c_{i})\Im(c_{j})\,.$
The other properties of scalar products are easily checked.
Therefore the completion of the space of polynomials on $\mathbb{R}$ equipped with the scalar product $\left\langle\cdot,\cdot\right\rangle_{s}$ and corresponding norm $\left\|\cdot\right\|_{s}^{2}:=\left\langle\cdot,\cdot\right\rangle_{s}$ gives us a Hilbert space that we call $\mathcal{H}_{s}$. We also note that the multiplication operator on the space of all polynomials over $\mathbb{R}$, $p(x)\xmapsto{T_{x}}xp(x)$ is a symmetric and densely defined operator on $\mathcal{H}_{s}$. Indeed,
$\left\langle xp(x),q(x)\right\rangle=\mathscr{L}(\overline{xp(x)}q(x))=\mathscr{L}(\overline{p(x)}xq(x))=\left\langle p(x),xq(x)\right\rangle\,.$
It is also known that $T_{x}$ has deficiency indices of either $(0,0)$ or $(1,1)$ [22, Corollary 6.7].
Theorem 4.4 (Theorem 6.10 [22]).
The moment problem for a positive definite sequence $s$ is determinate if and only if the multiplication operator $T_{x}$ on $\mathcal{H}_{s}$ has deficiency indices $(0,0)$, or in other words, $T_{x}$ is essentially self-adjoint on the space $\mathcal{H}_{s}$. If this holds and $\mu$ is the unique representing measure for $s$, then the space of polynomials on $\mathbb{R}$ is dense in the space $L^{2}(\mathbb{R},\mu)$, i.e., $\mathcal{H}_{s}\cong L^{2}(\mathbb{R},\mu)$.
Corollary 4.5 (Corollary 6.11 [22]).
Suppose that $\mu$ is a positive Radon measure on $\mathbb{R}$. Then the moment sequence $s$ defined by (4.1) is determinate if and only if the space of all polynomials on $\mathbb{R}$ is dense in $L^{2}(\mathbb{R},(1+x^{2})\mathrm{d}\mu)$.
Remark 4.6.
Under the condition that the sequence $s$ is positive definite, each representing measure $\mu$ for the moment problem on $\mathbb{R}$ has infinite support, that is, its support contains an infinite subset of $\mathbb{R}$ ([22, Proposition 3.11]). We note that for the case that $s$ is a positive semidefinite sequence that the determinacy of the Hamburger moment problem is under full control. Indeed, under the case that there exists a real polynomial $p(x)=\sum_{i=0}^{n}\xi_{i}x^{i}$, $\xi_{i}\in\mathbb{R}$ for all $i$, such that the moment functional applied to $p^{2}$ is zero, i.e., $0=\sum_{i,j=0}^{n}\xi_{i}\xi_{j}s_{i+j}=\mathscr{L}(p^{2})$, then any positive Radon measure representing $\mathscr{L}$ must have support at exactly finitely many points. In the case that a representing measure $\mu$ for $s$ is supported on finitely many points, and thus has compact support, it is the only representing measure for the sequence $s$, i.e., the Hamburger moment problem is determinate ([22, Corollary 4.2]).
In the following proposition we give a sufficient condition that is also practical to check, known as Carleman’s Condition, that ensures determinacy of the Hamburger moment problem. The statement and proof may be found in [22, Theorem 4.3].
Proposition 4.7 (Carleman’s Condition).
Suppose that $s\equiv(s_{n})_{n\in\mathbb{N}_{0}}$ is a positive semidefinite sequence. If $s$ satisfies the Carleman Condition
(4.5) $\sum_{n=0}^{\infty}s_{2n}^{-\frac{1}{2n}}=+\infty$
then the Hamburger moment problem is determinate.
4.2. Structural Properties of $\overline{\mathcal{K}\left(A,g\right)}$
We now begin our investigation of the structural properties of the Krylov subspace $\overline{\mathcal{K}\left(A,g\right)}$ under the condition that $A$ is a self-adjoint operator on a Hilbert space $\mathcal{H}$. In order to proceed, we first describe some important classes of vectors, namely the bounded, analytic, and quasi-analytic classes. More information on these classes may be found in [21, Chapter 7].
Definition 4.8.
Let $A$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$, and let $g\in C^{\infty}(A)$. We say that
  • (i)
    $g$ is of the bounded class with respect to $A$, i.e., $g\in\mathcal{D}^{b}(A)$, if there exists some $B_{g}>0$ such that $\left\|A^{n}g\right\|_{\mathcal{H}}\leqslant B_{g}^{n}$ for all $n\in\mathbb{N}$,
  • (ii)
    $g$ is of the analytic class with respect to $A$, i.e., $g\in\mathcal{D}^{a}(A)$, if there exists some $C_{g}>0$ such that $\left\|A^{n}g\right\|_{\mathcal{H}}\leqslant C_{g}^{n}n!$,
  • (iii)
    and $g$ is of the quasi-analytic class with respect to $A$, i.e., $g\in\mathcal{D}^{qa}(A)$, if
    $\sum_{n=1}^{\infty}\left\|A^{n}g\right\|_{\mathcal{H}}^{-\frac{1}{n}}=+\infty\,.$
Clearly we have that $\mathcal{D}^{b}(A)\subset\mathcal{D}^{a}(A)\subset\mathcal{D}^{qa}(A)$. Moreover, $\mathcal{D}^{b}(A)$ and $\mathcal{D}^{a}(A)$ are linear subspaces of $\mathcal{H}$, while $\mathcal{D}^{qa}(A)$ is not necessarily so.
We note that for $A$ a self-adjoint operator and $g\in C^{\infty}(A)$ we have a positive bounded Radon measure on the real line $\mu_{g}^{(A)}(\Omega):=\left\langle E^{(A)}(\Omega)g,g\right\rangle$, for $\Omega\subset\mathbb{R}$ a Borel measurable set and $E^{(A)}(\Omega)$ the unique projection valued spectral measure for $A$. Therefore, we can define a linear functional $\mathscr{L}_{\mu}$ on the space of polynomials as follows
(4.6) $\mathscr{L}_{\mu}(p(\lambda)):=\int_{\mathbb{R}}p(\lambda)\,\mathrm{d}\mu_{g}^{(A)}\,,$
and this gives rise to a positive semidefinite moment sequence $s\equiv(s_{n})_{n\in\mathbb{N}_{0}}$, where
(4.7) $s_{n}:=\mathscr{L}_{\mu}(\lambda^{n})\,,$
for all $n\in\mathbb{N}_{0}$. It is simple to see that $s$ is positive semidefinite: indeed as $\mu_{g}^{(A)}$ is a positive bounded measure and $g\in C^{\infty}(A)$, $\mathscr{L}_{\mu}(\lambda^{n})<+\infty$ for all $n\in\mathbb{N}_{0}$ and also $\mathscr{L}(p^{2}(\lambda))\geqslant 0$ for any real polynomial $p$. Therefore owing to Theorem 4.2 we have that $s$ is positive semidefinite. As the functional $\mathscr{L}_{\mu}$ is constructed from the positive Radon measure $\mu_{g}^{(A)}$ on $\mathbb{R}$ and $\lambda^{n}\in L^{1}(\mathbb{R},\mu_{g}^{(A)})$ we shall concern ourselves with the determinacy problem. In other words, we wish to know whether the measure $\mu_{g}^{(A)}$ is the only positive Radon measure that solves the Hamburger moment problem. To this end, we list below a sufficient condition for this to be the case.
Proposition 4.9.
Let $A$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$ and let $g\in\mathcal{D}^{qa}(A)$. Then the Hamburger moment problem with moment functional as in (4.6) is determinate. Moreover, $\overline{\mathcal{K}\left(A,g\right)}\cong L^{2}(\mathbb{R},\mu_{g}^{(A)})$.
Proof.
Let $\mu_{g}^{(A)}(\cdot)=\left\langle E^{(A)}(\cdot)g,g\right\rangle$ be the positive Radon measure generated by the projection valued spectral measure $E^{(A)}$ for $A$. By the functional calculus
$\left\|A^{n}g\right\|_{\mathcal{H}}^{2}=\int_{\mathbb{R}}\lambda^{2n}\,\mathrm{d}\mu_{g}^{(A)}\,,$
so that given the Hamburger moment sequence $s\equiv(s_{n})_{n\in\mathbb{N}_{0}}$ defined by $s_{n}:=\int_{\mathbb{R}}\lambda^{n}\,\mathrm{d}\mu_{g}^{(A)}$ for all $n\in\mathbb{N}_{0}$, we see that $s_{2n}=\left\|A^{n}g\right\|_{\mathcal{H}}^{2}$ for all $n\in\mathbb{N}_{0}$.
Then, as $g$ is quasi-analytic,
$\sum_{n=1}^{\infty}s_{2n}^{-\frac{1}{2n}}=\sum_{n=1}^{\infty}\left\|A^{n}g\right\|_{\mathcal{H}}^{-\frac{1}{n}}=+\infty\,,$
so that the Carleman Condition is satisfied, and therefore the Hamburger moment problem is determinate.
The proof that $\overline{\mathcal{K}\left(A,g\right)}\cong L^{2}(\mathbb{R},\mu_{g}^{(A)})$ is a direct consequence of Theorem 4.4–indeed we see that $\mathcal{H}_{s}=\overline{\mathcal{K}\left(A,g\right)}$. ∎
Remark 4.10.
We note that in the above Proposition, we actually have the more general result that $\overline{\mathcal{K}\left(A,g\right)}\cong L^{2}(\mathbb{R},\mu_{g}^{(A)})$ if and only if the Hamburger moment problem defined by (4.6) and (4.7) is determinate. As in the proof of Proposition 4.9 this is a direct consequence of Theorem 4.4.
We now use Corollary 4.5 in order to show that the Krylov core condition holds when $g$ is a vector that makes the Hamburger moment problem determinate.
Proposition 4.11.
Let $A$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$ and let $g\in C^{\infty}(A)$. If the Hamburger moment problem defined by (4.6) and (4.7) is determinate, then the Krylov core condition is satisfied, i.e., $\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A)=\overline{\mathcal{K}\left(A,g\right)}^{V}$. Moreover, we have the inclusions
(4.8) $A(\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A))\subset\overline{\mathcal{K}\left(A,g\right)}\,,\quad A(\mathcal{K}\left(A,g\right)^{\perp}\cap\mathcal{D}(A))\subset\mathcal{K}\left(A,g\right)^{\perp}\,.$
Proof.
It is already clear that $\overline{\mathcal{K}\left(A,g\right)}^{V}\subset\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A)$ so our proof will show that $\overline{\mathcal{K}\left(A,g\right)}^{V}\supset\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A)$. Indeed, take any vector $v\in\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A)$. Then the vector $v$ can be represented as a function $f$ in $L^{2}(\mathbb{R},\mu_{g}^{(A)})$ by way of the functional calculus ([4, Remark 7.7])
$v=\int_{\mathbb{R}}f(\lambda)\,\mathrm{d}E^{(A)}g\,.$
$\left\|Av\right\|_{\mathcal{H}}^{2}<+\infty$ as $v\in\mathcal{D}(A)$ and therefore
$\int_{\mathbb{R}}\lambda^{2}|f(\lambda)|^{2}\,\mathrm{d}\mu_{g}^{(A)}<+\infty\,,$
from which $f\in L^{2}(\mathbb{R},(1+\lambda^{2})\mu_{g}^{(A)})$.
As the Hamburger moment problem is determinate, by Corollary 4.5 we know that the polynomials are dense in $L^{2}(\mathbb{R},(1+\lambda^{2})\mu_{g}^{(A)})$ and therefore there exists a polynomial approximation $(p_{n})_{n\in\mathbb{N}}$ such that $p_{n}\xrightarrow{n\to\infty}f$ in $L^{2}(\mathbb{R},(1+\lambda^{2})\mu_{g}^{(A)})$. Thus
$\left\|p_{n}(A)g-f(A)g\right\|_{A}^{2}$ $=\left\|p_{n}(A)g-f(A)g\right\|_{\mathcal{H}}^{2}+\left\|Ap_{n}(A)g-Af(A)g\right\|_{\mathcal{H}}^{2}$
$=\int_{\mathbb{R}}(1+\lambda^{2})|p_{n}(\lambda)-f(\lambda)|^{2}\,\mathrm{d}\mu_{g}^{(A)}\xrightarrow{n\to\infty}0\,,$
and therefore $v=f(A)g\in\overline{\mathcal{K}\left(A,g\right)}^{V}$. By Propositions 3.2 and 3.4 we have the inclusions (4.8). ∎
Remark 4.12.
A proof that the class of bounded vectors satisfies the Krylov core condition is presented in [4, Theorem 7.1] and also in [6, Theorem 4.4]. These two proofs were done in an explicit way using a regularising sequence of vectors ([4, Theorem 7.1]) and using a purely measure-theoretic approach ([6, Theorem 4.4]). The proof of Proposition 4.11 is very direct and straightforward in comparison to the previous two studies, yet it requires the non-trivial knowledge of results in the theory of the moment problem in one dimension.
Remark 4.13.
At this point we remark on the two open questions posed in [4]. These were
  • (Q1)
    When $A$ is self-adjoint and $g\in C^{\infty}(A)$, is it true that $\overline{\mathcal{K}\left(A,g\right)}^{V}=\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A)$?
  • (Q2)
    When $A$ is self-adjoint and $g\in C^{\infty}(A)$ is it true that $A\overline{\mathcal{K}\left(A,g\right)}\cap\mathcal{D}(A)\subset\overline{\mathcal{K}\left(A,g\right)}$?
We note that in [4] a partial answer in the affirmative for both (Q1) and (Q2) was given in Theorem 7.1 therein, namely for the case in which $g\in\mathcal{D}^{b}(A)$. Our Proposition 4.11 expands on this answer by showing that the entire class of smooth vectors that result in a determinate Hamburger moment problem provide a positive answer to these two questions.
We shall now prove a stronger result than (4.8) for the class of bounded vectors, that will then be extended to larger classes by way of a perturbative analysis in Section 5. Before we do so, we shall state an important approximation theorem for continuous functions on $\mathbb{R}$ called Carleman’s Theorem (not to be confused with Proposition 4.7, namely Carleman’s Condition for the determinacy of the Hamburger moment problem).
Theorem 4.14 (Chapter IV, Section 3, Theorem 1 [13]).
Let $f:\mathbb{R}\to\mathbb{C}$ be a continuous function. Given any strictly positive function $\varepsilon:\mathbb{R}\to[0,\infty)$, there exists an entire function $h:\mathbb{C}\to\mathbb{C}$ such that
$|h(x)-f(x)|<\varepsilon(x)\,,\quad\forall\,x\in\mathbb{R}\,.$
Calculating the range of separations (Definition 2.8) between $\overline{\mathcal{K}\left(A,g\right)}$ and $A\mathcal{K}\left(A,g\right)^{\perp_{V}}$ gives us an indicator as to whether $\mathscr{I}_{\mathcal{K}}\left(A,g\right)=\{0\}$. Indeed, take any vector $v\in\mathcal{K}\left(A,g\right)^{\perp_{V}}$ and $u\in\mathcal{K}\left(A,g\right)$ such that $\left\|u\right\|_{\mathcal{H}}=1$ and $\left\|Av\right\|_{\mathcal{H}}=1$. Then
$\left\|u-Av\right\|_{\mathcal{H}}^{2}=2(1-\Re\left\langle u,Av\right\rangle)\,,$
and
$\left\langle u,Av\right\rangle=\int_{\mathbb{R}}p(\lambda)\lambda\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\,,$
where $p$ is a polynomial and $u=\int_{\mathbb{R}}p(\lambda)\,\mathrm{d}E^{(A)}(\lambda)g=p(A)g$. Our aim is to show that $\left\langle u,Av\right\rangle=0$ under certain circumstances. We now state and prove a proposition central to our structural investigations of the Krylov subspace structure that indeed shows that $\left\langle u,Av\right\rangle$ is zero when $g$ is of the bounded class of vectors.
Proposition 4.15.
Let $A$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$ and let $g\in\mathcal{D}^{b}(A)$. Then $\mathcal{K}\left(A,g\right)^{\perp_{V}}\subset\mathcal{K}\left(A,g\right)^{\perp}$ and $A\mathcal{K}\left(A,g\right)^{\perp_{V}}\subset\mathcal{K}\left(A,g\right)^{\perp}$.
Proof.
Given any $u\in\mathcal{K}\left(A,g\right)$ and any $v\in\mathcal{K}\left(A,g\right)^{\perp_{V}}$,
$0=\left\langle u,v\right\rangle_{V}=\int_{\mathbb{R}}p(\lambda)(1+\lambda^{2})\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\,,$
where $u=\int_{\mathbb{R}}p(\lambda)\,\mathrm{d}E^{(A)}g$ for a polynomial $p$. Owing to the invariance of $\mathcal{K}\left(A,g\right)$ under $A$, we also have that
$0=\left\langle A^{k}u,v\right\rangle_{V}=\left\langle A^{k}p(A)g,v\right\rangle_{V}=\int_{\mathbb{R}}p(\lambda)(1+\lambda^{2})\lambda^{k}\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\,,$
for any polynomial $p$ and for all $k\in\mathbb{N}_{0}$. We will construct a sequence of polynomials $(p_{m})_{m\in\mathbb{N}}$ such that we have the limit
(*) $0=\left\langle A^{k}p_{m}(A)g,v\right\rangle_{V}\xrightarrow{m\to\infty}\left\langle A^{k}g,v\right\rangle\,,$
which is tantamount to the convergence of the integrals
(**) $\int_{\mathbb{R}}p_{m}(\lambda)(1+\lambda^{2})\lambda^{k}\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\xrightarrow{m\to\infty}\int_{\mathbb{R}}\lambda^{k}\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\,,$
which then implies that $A^{k}g\perp v$, for $k=0,1$. Then for all $k\in\mathbb{N}_{0}$ we will show that $\left\langle A^{k}g,v\right\rangle=0$.
To achieve this, we invoke Carleman’s Theorem (Theorem 4.14), noting that $(1+\lambda^{2})^{-1}$ is a continuous function on $\mathbb{R}$. We choose an $\varepsilon>0$ and let the positive error function be $\varepsilon(\lambda)=\varepsilon$. Therefore there exists an entire function $h_{\varepsilon}:\mathbb{C}\to\mathbb{C}$ such that $|h_{\varepsilon}(\lambda)-(1+\lambda^{2})^{-1}|<\varepsilon$ for all $\lambda\in\mathbb{R}$. Let us define a polynomial sequence as follows
$q_{n}^{(h_{\varepsilon})}(\lambda):=\sum_{i=0}^{n}\frac{h_{\varepsilon}^{(i)}(0)}{i!}\lambda^{i}\,,\quad n\in\mathbb{N}_{0}\,,$
which is the $n$-th order truncation of the Maclaurin series for $h_{\varepsilon}(\lambda)$. As $h_{\varepsilon}$ is entire, the series has an infinite radius of convergence, so $q_{n}^{(h_{\varepsilon})}\xrightarrow{n\to\infty}h$ uniformly on compact subsets of $\mathbb{C}$. The function
$f(\lambda):=\sum_{i=0}^{\infty}\frac{|h_{\varepsilon}^{(i)}(0)|}{i!}|\lambda|^{i}$
dominates both $h_{\varepsilon}$ and each $q_{n}^{(h_{\varepsilon})}$, so that $2f(\lambda)(1+\lambda^{2})|\lambda|^{k}$ dominates $|(q_{n}^{(h_{\varepsilon})}(\lambda)-h_{\varepsilon}(\lambda))(1+\lambda^{2})\lambda^{k}|$ for all $k\in\mathbb{N}_{0}$. (Note that while $f$ depends on $h_{\varepsilon}$, we leave out the dependence in its notation in order to avoid obfuscation.)
To show that $f(\lambda)$ actually exists for each $\lambda\in\mathbb{R}$ we consider the Cauchy estimate for $h_{\varepsilon}^{(i)}(0)$, namely
$|h_{\varepsilon}^{(i)}(0)|\leqslant\frac{M_{R}i!}{R^{i}}\,,$
where $R>0$ and $M_{R}=\max_{z=R}|h_{\varepsilon}(z)|$. Then for fixed $\lambda\in\mathbb{R}$
$\frac{|h_{\varepsilon}^{(i)}(0)|}{i!}|\lambda|^{i}\leqslant M_{R}\frac{|\lambda|^{i}}{R^{i}}\,,$
and thus the summation converges as we may choose $R>|\lambda|$ without restriction.
We claim that $f(\lambda)(1+\lambda^{2})|\lambda|^{k}\in L^{2}(\mathbb{R},\mu_{g}^{(A)})$ for all $k\in\mathbb{N}_{0}$. Indeed, consider $f^{2}(\lambda)$
$|f(\lambda)|^{2}=f^{2}(\lambda)=\sum_{i=0}^{\infty}\frac{|h_{\varepsilon}^{(i)}(0)|}{i!}|\lambda|^{i}\sum_{j=0}^{\infty}\frac{|h_{\varepsilon}^{(j)}(0)|}{j!}|\lambda|^{j}\,,$
so that owing to the (absolute) convergence of both series at each $\lambda\in\mathbb{R}$ we can replace the double summation with the Cauchy product
$f^{2}(\lambda)=\sum_{i=0}^{\infty}c_{i}|\lambda|^{i}\,,\quad c_{i}=\sum_{l=0}^{i}\frac{|h_{\varepsilon}^{(l)}(0)|}{l!}\frac{|h_{\varepsilon}^{(i-l)}(0)|}{(i-l)!}\,.$
Taking the Cauchy estimate for some $R>0$ (to be chosen later), we see that
$c_{i}\leqslant M_{R}^{2}\sum_{l=0}^{i}\frac{1}{R^{i}}=\frac{(i+1)M_{R}^{2}}{R^{i}}\,,$
and therefore
$f^{2}(\lambda)\leqslant M_{R}^{2}\sum_{i=0}^{\infty}\frac{i+1}{R^{i}}|\lambda|^{i}\,.$
Indeed we see that the right side converges for each $\lambda\in\mathbb{R}$ by choosing some $R$ large enough e.g., for fixed $\lambda$ one may choose $R>4|\lambda|$ noting that $i+1<2^{i}$ for all $i\in\mathbb{N}$ and $2|\lambda|\geqslant|\lambda|$ for all $\lambda\in\mathbb{R}$.
Putting this into the spectral integral, we have by the Monotone Convergence Theorem and the Hölder inequality
$\int_{\mathbb{R}}f^{2}(\lambda)(1+\lambda^{2})^{2}\lambda^{2k}\,\mathrm{d}\mu_{g}^{(A)}$ $\leqslant M_{R}^{2}\sum_{i=0}^{\infty}\int_{\mathbb{R}}\frac{i+1}{R^{i}}|\lambda|^{i}(1+\lambda^{2})^{2}\lambda^{2k}\,\mathrm{d}\mu_{g}^{(A)}$
$\leqslant M_{R}^{2}\sum_{i=0}^{\infty}\frac{i+1}{R^{i}}\left(\int_{\mathbb{R}}\lambda^{2i}\,\mathrm{d}\mu_{g}^{(A)}\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}(1+\lambda^{2})^{4}\lambda^{4k}\,\mathrm{d}\mu_{g}^{(A)}\right)^{\frac{1}{2}}$
$\leqslant\tilde{\mathscr{C}}_{k}M_{R}^{2}\sum_{i=0}^{\infty}\frac{i+1}{R^{i}}B_{g}^{i}\,,$
where $\tilde{\mathscr{C}}_{k}=\left(\int_{\mathbb{R}}(1+\lambda^{2})^{4}\lambda^{4k}\,\mathrm{d}\mu_{g}^{(A)}\right)^{\frac{1}{2}}<\infty$ for all $k\in\mathbb{N}_{0}$, so that the series on the right side of the last inequality converges for any choice of $R>4\max\{1,B_{g}\}$. This proves that $f(\lambda)(1+\lambda^{2})|\lambda|^{k}\in L^{2}(\mathbb{R},\mu_{g}^{(A)})$ for all $k\in\mathbb{N}_{0}$.
Thus considering $k=0,1$
(a) $\begin{split}\bigg{|}\int_{\mathbb{R}}(q^{(h_{\varepsilon})}_{n}(\lambda)-h_{\varepsilon}(\lambda))(1+&\lambda^{2})\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\bigg{|}^{2}\\ &\leqslant\int_{\mathbb{R}}|q^{(h_{\varepsilon})}_{n}(\lambda)-h_{\varepsilon}(\lambda)|^{2}(1+\lambda^{2})^{2}\,\mathrm{d}\mu_{g}^{(A)}\int_{\mathbb{R}}\,\mathrm{d}\mu_{v}^{(A)}\\ &=\left\|v\right\|_{\mathcal{H}}^{2}\int_{\mathbb{R}}|q_{n}^{(h_{\varepsilon})}(\lambda)-h_{\varepsilon}(\lambda)|^{2}(1+\lambda^{2})^{2}\,\mathrm{d}\mu_{g}^{(A)}\,,\end{split}$
and
(b) $\begin{split}\bigg{|}\int_{\mathbb{R}}(q_{n}^{(h_{\varepsilon})}(\lambda)-h_{\varepsilon}(\lambda))&(1+\lambda^{2})\lambda\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\bigg{|}^{2}\\ &\leqslant\int_{\mathbb{R}}|q_{n}^{(h_{\varepsilon})}(\lambda)-h_{\varepsilon}(\lambda)|^{2}(1+\lambda^{2})^{2}\,\mathrm{d}\mu_{g}^{(A)}\int_{\mathbb{R}}\lambda^{2}\,\mathrm{d}\mu_{v}^{(A)}\\ &=\left\|Av\right\|_{\mathcal{H}}^{2}\int_{\mathbb{R}}|q_{n}^{(h_{\varepsilon})}(\lambda)-h_{\varepsilon}(\lambda)|^{2}(1+\lambda^{2})^{2}\,\mathrm{d}\mu_{g}^{(A)}\,,\end{split}$
where we have used [21, Lemma 4.8(ii)]. Therefore, the convergence of $(q_{n}^{(h_{\varepsilon})}-h_{\varepsilon})(1+\lambda^{2})\xrightarrow{n\to\infty}0$ in $L^{2}(\mathbb{R},\mu_{g}^{(A)})$ implies convergence to zero of the left sides of the inequalities (a) and (b).
As $|(q_{n}^{(h_{\varepsilon})}(\lambda)-h_{\varepsilon}(\lambda))(1+\lambda^{2})|\leqslant 2f(\lambda)(1+\lambda^{2})\in L^{2}(\mathbb{R},\mu_{g}^{(A)})$ and $q_{n}^{(h_{\varepsilon})}(\lambda)(1+\lambda^{2})\xrightarrow{n\to\infty}h_{\varepsilon}(\lambda)(1+\lambda^{2})$ pointwise, Lebesgue Dominated Convergence gives
$\int_{\mathbb{R}}|q_{n}^{(h_{\varepsilon})}(\lambda)-h_{\varepsilon}(\lambda)|^{2}(1+\lambda^{2})^{2}\,\mathrm{d}\mu_{g}^{(A)}\xrightarrow{n\to\infty}0\,,$
and therefore we have the quantities on the left sides of (a) and (b) converge to 00 as $n\to\infty$.
For $k=0,1$ there exists an $n_{\varepsilon}\in\mathbb{N}$ such that
$\left|\int_{\mathbb{R}}q_{n_{\varepsilon}}^{(h_{\varepsilon})}(\lambda)(1+\lambda^{2})\lambda^{k}\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle-\int_{\mathbb{R}}h_{\varepsilon}(\lambda)(1+\lambda^{2})\lambda^{k}\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\right|<\varepsilon\,.$
Therefore, for $k=0,1$,
$\bigg{|}\int_{\mathbb{R}}\bigg{(}q_{n_{\varepsilon}}^{(h_{\varepsilon})}(\lambda)$ $-\frac{1}{1+\lambda^{2}}\bigg{)}(1+\lambda^{2})\lambda^{k}\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\bigg{|}$
$\leqslant\left|\int_{\mathbb{R}}(q_{n_{\varepsilon}}^{(h_{\varepsilon})}(\lambda)-h_{\varepsilon}(\lambda))(1+\lambda^{2})\lambda^{k}\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\right|$
$\qquad+\left|\int_{\mathbb{R}}\left(h_{\varepsilon}(\lambda)-\frac{1}{1+\lambda^{2}}\right)(1+\lambda^{2})\lambda^{k}\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle\right|$
$<\varepsilon+\int_{\mathbb{R}}\left|h_{\varepsilon}(\lambda)-\frac{1}{1+\lambda^{2}}\right|(1+\lambda^{2})|\lambda^{k}|\,\mathrm{d}\left\langle E^{(A)}(\lambda)g,v\right\rangle$
$\leqslant\varepsilon+\left(\int_{\mathbb{R}}\left|h_{\varepsilon}(\lambda)-\frac{1}{1+\lambda^{2}}\right|^{2}(1+\lambda^{2})^{2}\,\mathrm{d}\mu_{g}^{(A)}\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\lambda^{2k}\,\mathrm{d}\mu_{v}^{(A)}\right)^{\frac{1}{2}}$
$\leqslant\varepsilon+\mathscr{C}_{k}\varepsilon\left\|(\mathbbm{1}+A^{2})g\right\|_{\mathcal{H}}$
where $\mathscr{C}_{0}=\left\|v\right\|_{\mathcal{H}}$ and $\mathscr{C}_{1}=\left\|Av\right\|_{\mathcal{H}}$, and we have used [21, Lemma 4.8(ii)].
We may set $\varepsilon=\frac{1}{m}$ for $m\in\mathbb{N}$, and extract a sequence $(p_{m})_{m\in\mathbb{N}}$ with the polynomials $p_{m}=q_{n_{1/m}}^{(h_{1/m})}$ for all $m\in\mathbb{N}$. We note that in the above construction, the choice of $(p_{m})_{m\in\mathbb{N}}$ is independent of the choice of vector $v\in\mathcal{K}\left(A,g\right)^{\perp_{V}}$, and therefore (**4.2), and consequently (*4.2), are satisfied for $k=0,1$.
We now turn our attention to the case for which $k>1$. The consequence of the limit in (*4.2) for $k=0,1$ is that
(***) $\left\langle g,v\right\rangle=0\,\quad\text{and}\quad 0=\left\langle Ag,v\right\rangle=\left\langle g,Av\right\rangle\,.$
Applying the fact that $v\in\mathcal{K}\left(A,g\right)^{\perp_{V}}$, we have
$0=\left\langle g,v\right\rangle_{V}\implies\left\langle g,v\right\rangle=-\left\langle Ag,Av\right\rangle=-\left\langle A^{2}g,v\right\rangle\,,$
and
$0=\left\langle Ag,v\right\rangle_{V}\implies\left\langle Ag,v\right\rangle=-\left\langle A^{2}g,Av\right\rangle=-\left\langle A^{3}g,v\right\rangle\,,$
so that $\left\langle A^{2}g,v\right\rangle=\left\langle A^{3}g,v\right\rangle=0$ from (***4.2). Then,
$0=\left\langle A^{2}g,v\right\rangle_{V}\implies 0=\left\langle A^{2}g,v\right\rangle=-\left\langle A^{3}g,Av\right\rangle=-\left\langle A^{4}g,v\right\rangle\,,$
and
$0=\left\langle A^{3}g,v\right\rangle_{V}\implies 0=\left\langle A^{3}g,v\right\rangle=-\left\langle A^{4}g,Av\right\rangle=-\left\langle A^{5}g,v\right\rangle\,.$
Continuing in this way, we eventually get that $\left\langle A^{k}g,v\right\rangle=0$ and $\left\langle A^{k}g,Av\right\rangle=0$ for all $k\in\mathbb{N}_{0}$.
From the arbitrary choice for $v$ we see that both $\mathcal{K}\left(A,g\right)^{\perp_{V}}\subset\mathcal{K}\left(A,g\right)^{\perp}$ and $A\mathcal{K}\left(A,g\right)^{\perp_{V}}\subset\mathcal{K}\left(A,g\right)^{\perp}$. ∎
As such, for any vector $u\in\overline{\mathcal{K}\left(A,g\right)}$ such that $\left\|u\right\|_{\mathcal{H}}=1$ and any vector $v\in\mathcal{K}\left(A,g\right)^{\perp_{V}}$ such that $\left\|Av\right\|_{\mathcal{H}}=1$, we have
$\left\|u-Av\right\|_{\mathcal{H}}^{2}=2(1-\Re\left\langle u,Av\right\rangle)=2$
so that $\mathcal{S}(\mathcal{K}\left(A,g\right),A\mathcal{K}\left(A,g\right)^{\perp_{V}})=\{\sqrt{2}\}$ and therefore, from Proposition 2.10 we have that $\mathscr{I}_{\mathcal{K}}\left(A,g\right)=\{0\}$. This is the statement of the next theorem.
Theorem 4.16.
Let $A:\mathcal{H}\to\mathcal{H}$ be a self-adjoint operator, and $g\in\mathcal{H}$ be a bounded vector with respect to $A$. Then $\mathscr{I}_{\mathcal{K}}\left(A,g\right)=\{0\}$. Moreover, if $g\in\mathrm{ran}A$ then the inverse linear problem $Af=g$ has a solution $f\in\overline{\mathcal{K}\left(A,g\right)}^{V}$.
Proof.
The Krylov solvability is a consequence of Theorem 2.6. ∎
5. Perturbative Analysis for Unbounded Self-Adjoint Operators and Krylov Inner Approximations
We are now left with the question: what happens to the Krylov subspace structure when one considers $g$ as a more general vector (e.g., analytic as opposed to bounded)? Specifically, we wish to know what happens to the Krylov intersection (2.2) for more general vectors, and we shall investigate this by using the bounded class of vectors in a perturbation analysis.
Lemma 5.1.
Let $A:\mathcal{H}\to\mathcal{H}$ be a self-adjoint linear operator. Then the class of bounded vectors $\mathcal{D}^{b}(A)$ forms a core of $A$, and moreover given any $f\in\mathcal{D}(A)$ the sequence of bounded vectors $(\chi_{[-n,n]}(A)f)_{n\in\mathbb{N}}$ converges in graph norm to $f$.
Proof.
Let $f\in\mathcal{D}(A)$, and consider the sequence of vectors $(\chi_{[-n,n]}(A)f)_{n\in\mathbb{N}}$. We clearly know that each of these vectors $f_{n}:=\chi_{[-n,n]}(A)f$ is a bounded vector as
$\left\|A^{k}f_{n}\right\|_{\mathcal{H}}^{2}=\int_{\mathbb{R}}\lambda^{2k}\chi_{[-n,n]}(\lambda)\,\mathrm{d}\mu_{f}^{(A)}\leqslant n^{2k}\left\|f\right\|_{\mathcal{H}}^{2}\,.$
We also show that $f_{n}\xrightarrow{\left\|\cdot\right\|_{A}}f$ as $n\to\infty$. Indeed,
$\left\|f-f_{n}\right\|_{A}^{2}=\int_{\mathbb{R}}(1+\lambda^{2})|1-\chi_{[-n,n]}(\lambda)|^{2}\,\mathrm{d}\mu_{f}^{(A)}\xrightarrow{n\to\infty}0\,,$
from the Lebesgue Dominated Convergence theorem. ∎
Lemma 5.2.
Let $A:\mathcal{H}\to\mathcal{H}$ be a self-adjoint operator on Hilbert space $\mathcal{H}$, let $f\in\mathbb{C}^{\infty}(A)$, and consider the sequence of vectors $(f_{n})_{n\in\mathbb{N}}$ defined by $f_{n}=\chi_{[-n,n]}(A)f$ for all $n\in\mathbb{N}$. Then for any polynomial $p$, we have that $p(A)f_{n}\xrightarrow{\left\|\cdot\right\|_{A}}p(A)f$. Moreover, $\left\|p(A)f_{n}\right\|_{\mathcal{H}}\leqslant\left\|p(A)f\right\|_{\mathcal{H}}$ and $\left\|p(A)f_{n}\right\|_{A}\leqslant\left\|p(A)f\right\|_{A}$.
Proof.
$\left\|p(A)(f-f_{n})\right\|_{A}^{2}=\int_{\mathbb{R}}(1+\lambda^{2})|p(\lambda)|^{2}|1-\chi_{[-n,n]}(\lambda)|^{2}\,\mathrm{d}\mu_{f}^{(A)}\,.$
The integral vanishes as $n\to\infty$ due to the Lebesgue Dominated Convergence Theorem. Indeed $(1+\lambda^{2})|p(\lambda)|^{2}$ is a suitable dominating function and the integrand vanishes pointwise. The final inequality is obvious from spectral integrals and the relation $\chi_{[-n,n]}(\lambda)\leqslant 1$ for all $\lambda\in\mathbb{R}$. ∎
Given $A$ a self-adjoint operator on a Hilbert space $\mathcal{H}$ and $g\in C^{\infty}(A)$ we define the following subspace
(5.1) $\mathfrak{L}:=\{h(A)g\,|\,h\in L^{2}(\mathbb{R},\mu_{g}^{(A)})\}\,.$
It is obvious that $\mathfrak{L}\cong L^{2}(\mathbb{R},\mu_{g}^{(A)})$ and also that $\overline{\mathcal{K}\left(A,g\right)}\subset\mathfrak{L}$. Moreover, we have that $\mathfrak{L}\cap\mathcal{D}(A)\cong L^{2}(\mathbb{R},(1+\lambda^{2})\mu_{g}^{(A)})$ in the space $V$, i.e., there exists an isometric isomorphism between $L^{2}(\mathbb{R},(1+\lambda^{2})\mu_{g}^{(A)})$ and $\mathfrak{L}\cap\mathcal{D}(A)$ in the Hilbert space $\mathcal{D}(\mathcal{H})$ equipped with the graph scalar product $\left\langle\cdot,\cdot\right\rangle_{V}$ and graph norm $\left\|\cdot\right\|_{A}$. Indeed $h\in L^{2}(\mathbb{R},(1+\lambda^{2})\mu_{g}^{(A)})$ if and only if $h(A)g\in\mathcal{D}(A)$ which is equivalent to the requirement $h(A)g\in\mathfrak{L}\cap\mathcal{D}(A)$. According to Theorem 4.4, $\mathfrak{L}=\overline{\mathcal{K}\left(A,g\right)}$ if and only if the Hamburger moment problem generated by (4.6) and (4.7) is determinate.
We now demonstrate appropriate limiting sequences for approximating the subspaces $\mathfrak{L}$ and $\overline{\mathcal{K}\left(A,g\right)}$ in terms of the weak-gap metric $\hat{d}_{w}$ from [7] when we have that $\mathcal{H}$ is separable.
Proposition 5.3.
Let $A$ be self-adjoint on a separable Hilbert space $\mathcal{H}$, and let $g\in C^{\infty}(A)$. Then there exists a sequence of bounded vectors $(g_{n})_{n\in\mathbb{N}}$ such that $g_{n}\xrightarrow{\left\|\cdot\right\|_{A}}g$, and moreover
  • (i)
    $\overline{\mathcal{K}\left(A,g_{n}\right)}\subset\overline{\mathcal{K}\left(A,g_{n+1}\right)}\subset\mathfrak{L}$ for all $n$,
  • (ii)
    and $\overline{\mathcal{K}\left(A,g_{n}\right)}\xrightarrow{\hat{d}_{w}}\mathfrak{L}$, where $\hat{d}_{w}$ is the weak-gap metric based on the weak topology of $B_{\mathcal{H}}$.
Proof.
We begin by proving part (i). Let $g_{n}:=\chi_{[-n,n]}(A)g=\int_{\mathbb{R}}\chi_{[-n,n]}(\lambda)\,\mathrm{d}E^{(A)}(\lambda)g$, where $E^{(A)}$ is the unique projection valued spectral measure for the self-adjoint operator $A$. By Lemma 5.1 we know that $g_{n}\xrightarrow{\left\|\cdot\right\|_{A}}g$.
We show that $\overline{\mathcal{K}\left(A,g_{n}\right)}\subset\overline{\mathcal{K}\left(A,g_{n+1}\right)}$, and $\overline{\mathcal{K}\left(A,g_{n}\right)}\subset\mathfrak{L}$ for all $n\in\mathbb{N}$. As $g_{n}\in\mathcal{D}^{b}(A)$, $\overline{\mathcal{K}\left(A,g_{n}\right)}\cong L^{2}(\mathbb{R},\mu_{g_{n}}^{(A)})$ for all $n\in\mathbb{N}_{0}$. So, given any $h\in L^{2}(\mathbb{R},\mu_{g_{n}}^{(A)})$
$\int_{\mathbb{R}}|h(\lambda)|^{2}\,\mathrm{d}\mu_{g_{n}}^{(A)}$ $=\int_{\mathbb{R}}|h(\lambda)|^{2}\chi_{[-n,n]}(\lambda)\,\mathrm{d}\mu_{g}^{(A)}$
$=\int_{\mathbb{R}}|h(\lambda)|^{2}\chi_{[-n,n]}(\lambda)\chi_{[-n-1,n+1]}(\lambda)\,\mathrm{d}\mu_{g}^{(A)}$
$=\int_{\mathbb{R}}|h(\lambda)|^{2}\chi_{[-n,n]}(\lambda)\,\mathrm{d}\mu_{g_{n+1}}^{(A)}$
so that $h(\lambda)\chi_{[-n,n]}(\lambda)\in L^{2}(\mathbb{R},\mu_{g_{n+1}}^{(A)})$. We see that
$\overline{\mathcal{K}\left(A,g_{n+1}\right)}\ni h(A)\chi_{[-n,n]}(A)g_{n+1}=h(A)\chi_{[-n,n]}(A)\chi_{[-n-1,n+1]}(A)g=h(A)g_{n}\,.$
We also have that $h(A)g_{n}=h(A)\chi_{[-n,n]}(A)g$ implies $h(\lambda)\chi_{[-n,n]}(\lambda)\in L^{2}(\mathbb{R},\mu_{g}^{(A)})$ and thus $h(A)g_{n}\in\mathfrak{L}$.
We now prove part (ii). By Lemma 3.9 we know that there exists some $M\subset B_{\mathcal{H}}$ such that $\mathcal{K}_{n}\xrightarrow{\hat{d}_{w}}M$, where we use the shorthand $\mathcal{K}_{n}:=\overline{\mathcal{K}\left(A,g_{n}\right)}\cap B_{\mathcal{H}}$ and $\mathcal{L}:=\mathfrak{L}\cap B_{\mathcal{H}}$.
We claim that $M=\mathcal{L}$. Indeed, from Theorem 3.5(iii), we have
$M=\{u\in B_{\mathcal{H}}\,|\,u_{n}\rightharpoonup u\text{ for a sequence }(u_{n})_{n\in\mathbb{N}}\text{ with }u_{n}\in M_{n}\}\,.$
Let $u\in\mathcal{L}$ be given, so there exists some $f\in L^{2}(\mathbb{R},\mu_{g}^{(A)})$ such that $f(A)g=u$. Let $f_{n}(\lambda):=f(\lambda)\chi_{[-n,n]}(\lambda)$, so that $f_{n}(A)g=f_{n}(A)g_{n}\in L^{2}(\mathbb{R},\mu_{g_{n}}^{(A)})$ which implies $f_{n}(A)g\in\mathcal{K}_{n}$ as $\left\|f_{n}(A)g\right\|_{\mathcal{H}}\leqslant\left\|f(A)g\right\|_{\mathcal{H}}\leqslant 1$. Letting $u_{n}:=f_{n}(A)g$,
$\left\|u_{n}-u\right\|_{\mathcal{H}}^{2}=\int_{\mathbb{R}}|f_{n}(\lambda)-f(\lambda)|^{2}\,\mathrm{d}\mu_{g}^{(A)}\xrightarrow{n\to\infty}0\,,$
from the Lebesgue Dominated Convergence Theorem. Therefore, there exists a sequence $(u_{n})_{n\in\mathbb{N}}$ such that $u_{n}\in\mathcal{K}_{n}$ for all $n\in\mathbb{N}$ and $u_{n}\to u$, from which $u\in M$. This gives us the inclusion $\mathcal{L}\subset M$.
It is clear that $M\subset\mathcal{L}$ given that $\mathcal{K}_{n}\subset\mathcal{L}$ as proven above. Thus $M=\mathcal{L}$, and this completes the proof. ∎
Remark 5.4.
This proposition shows us that when $\mathfrak{L}=\overline{\mathcal{K}\left(A,g\right)}$ the Krylov subspace can be “inner approximated” by a sequence of nested Krylov subspaces. Indeed, it was shown in [7] that there always exists the property of inner approximability of Krylov subspaces when $A$ is a bounded linear operator.
More interestingly, it also shows us that by using a nested sequence of Krylov subspaces we can actually approximate the full space $\mathfrak{L}$ which could strictly contain the closed Krylov subspace, precisely the case when the Hamburger moment problem (4.6) and (4.7) is not determinate. This suggests that such an approximation scheme of Krylov subspaces could be advantageous in constructing or approximating vectors that would not be present in $\overline{\mathcal{K}\left(A,g\right)}$ when the Hamburger moment problem is indeterminate.
Corollary 5.5.
Consider $A$, $g$, and $\mathcal{H}$ as in the statement of Proposition 5.3 with the corresponding sequence $(g_{n})_{n\in\mathbb{N}}$ from the same proposition. Then $\mathcal{K}\left(A,g_{n}\right)^{\perp}\xrightarrow{\hat{d}_{w}}\mathfrak{L}^{\perp}$.
Proof.
This is a result of combining Propositions 5.3 and 3.11. ∎
Remark 5.6.
At this point we comment on the action of the orthogonal projection operator $P_{\mathcal{K}_{n}}:\mathcal{H}\to\mathcal{H}$ onto $\overline{\mathcal{K}\left(A,g_{n}\right)}$ from Proposition 5.3 and the proof therein for the choice of vectors $g_{n}:=\chi_{[-n,n]}(A)g$ for all $n\in\mathbb{N}$.
For any vector $v\in\mathfrak{L}$ it turns out that we have an explicit formula for $P_{\mathcal{K}_{n}}v$ that can be written in terms of the functional calculus. As there is a corresponding $h\in L^{2}(\mathbb{R},\mu_{g}^{(A)})$ such that $h(A)g=v$ we have
(5.2) $P_{\mathcal{K}_{n}}v=\int_{\mathbb{R}}h(\lambda)\chi_{[-n,n]}(\lambda)\,\mathrm{d}E^{(A)}g=h(A)g_{n}\,.$
It is very clear that $P_{\mathcal{K}_{n}}v\xrightarrow{n\to\infty}v\in\mathfrak{L}$ by a Lebesgue Dominated Convergence argument, in agreement with Proposition 3.13. Indeed, the projection $P_{\mathcal{K}_{n}}v$ vector is the unique argument that minimises the square distance $\left\|u-v\right\|_{\mathcal{H}}^{2}$ over all $u\in\overline{\mathcal{K}\left(A,g_{n}\right)}$, so that
$\left\|\tilde{h}(A)g_{n}-h(A)g\right\|_{\mathcal{H}}^{2}$ $=\left\|\tilde{h}(A)\chi_{[-n,n]}(A)g-h(A)g\right\|_{\mathcal{H}}^{2}$
$=\int_{\mathbb{R}}|\tilde{h}(\lambda)\chi_{[-n,n]}(\lambda)-h(\lambda)|^{2}\,\mathrm{d}\mu_{g}^{(A)}$
$=\int_{[-n,n]}|\tilde{h}(\lambda)-h(\lambda)|^{2}\,\mathrm{d}\mu_{g}^{(A)}+\int_{\mathbb{R}\setminus[-n,n]}|h(\lambda)|^{2}\,\mathrm{d}\mu_{g}^{(A)}\,,$
where $\tilde{h}\in L^{2}(\mathbb{R},\mu_{g_{n}}^{(A)})$ is the representation of any vector $u\in\overline{\mathcal{K}\left(A,g_{n}\right)}\cong L^{2}(\mathbb{R},\mu_{g_{n}}^{(A)})$. The minimiser for the above integral is $\tilde{h}(\lambda)=\chi_{[-n,n]}(\lambda)h(\lambda)$ $\mu_{g}^{(A)}$-a.e., i.e., $\tilde{h}(\lambda)=h(\lambda)$ for $\mu_{g}^{(A)}$-a.e. $\lambda\in[-n,n]$. This implies $P_{\mathcal{K}_{n}}v=h(A)g_{n}=\chi_{[-n,n]}(A)v$.
As $V$ is a Hilbert space and separable when $\mathcal{H}$ is separable, we also have the notion of the weak-gap based on the weak topology of $B_{V}$ in $V$, namely $\hat{d}^{V}_{w}$. We shall exploit this construction for our next theorem.
Theorem 5.7.
Let $A$ be a self-adjoint operator on a separable Hilbert space $\mathcal{H}$, and let $g\in C^{\infty}(A)$ be such that the Hamburger moment problem (4.6) and (4.7) is determinate (for example, $g\in\mathcal{D}^{qa}(A)$).
  • (i)
    $\mathscr{I}_{\mathcal{K}}\left(A,g\right)=\{0\}$. Moreover, if $g\in\mathrm{ran}A$ then there exists a solution $f$ to $Af=g$ in $\overline{\mathcal{K}\left(A,g\right)}^{V}$.
  • (ii)
    For the sequence of vectors $(g_{n})_{n\in\mathbb{N}}$ as in the statement of Proposition 5.3, we have that $\overline{\mathcal{K}\left(A,g_{n}\right)}^{V}\xrightarrow{\hat{d}^{V}_{w}}\overline{\mathcal{K}\left(A,g\right)}^{V}$ and $\overline{\mathcal{K}\left(A,g_{n}\right)}^{\perp_{V}}\xrightarrow{\hat{d}^{V}_{w}}\overline{\mathcal{K}\left(A,g\right)}^{\perp_{V}}$, where $\hat{d}^{V}_{w}$ is the weak gap metric based on the unit ball $B_{V}=\{v\in V\,|\,\left\|v\right\|_{A}\leqslant 1\}$ in the Hilbert space $V$.
Proof.
We begin by proving statement (ii). Consider the sequence of bounded vectors $(g_{n})_{n\in\mathbb{N}}$ corresponding to the same sequence of vectors in the statement of Proposition 5.3. It is known that $\overline{\mathcal{K}\left(A,g\right)}\cong L^{2}(\mathbb{R},\mu_{g}^{(A)})$, i.e., $\mathfrak{L}=\overline{\mathcal{K}\left(A,g\right)}$, as the Hamburger moment problem is determinate (Theorem 4.4).
We already know that $\overline{\mathcal{K}\left(A,g_{n}\right)}$ and $\overline{\mathcal{K}\left(A,g\right)}$ satisfy the Krylov core condition (Proposition 4.11) and $\overline{\mathcal{K}\left(A,g_{n}\right)}\subset\overline{\mathcal{K}\left(A,g\right)}$ (Proposition 5.3), so that we end up with the inclusions
$\overline{\mathcal{K}\left(A,g_{n}\right)}^{V}\subset\overline{\mathcal{K}\left(A,g_{n+1}\right)}^{V}\subset\overline{\mathcal{K}\left(A,g\right)}^{V}\,,\quad\forall\,n\in\mathbb{N}\,,$
from which
$\mathcal{K}\left(A,g\right)^{\perp_{V}}\subset\mathcal{K}\left(A,g_{n+1}\right)^{\perp_{V}}\subset\mathcal{K}\left(A,g_{n}\right)^{\perp_{V}}\,,\quad\forall\,n\in\mathbb{N}\,.$
The nested structure of the Krylov subspaces means that $\overline{\mathcal{K}\left(A,g_{n}\right)}^{V}\cap B_{V}\xrightarrow{\hat{d}^{V}_{w}}C$ where $C$ is a weakly closed subset of $B_{V}$ (Lemma 3.9). Owing to the set inclusions above, we see that $C\subset\overline{\mathcal{K}\left(A,g\right)}^{V}\cap B_{V}$. Let $p$ be any polynomial, and by Lemma 5.2 we have
$\frac{\alpha p(A)g_{n}}{\left\|p(A)g\right\|_{A}}\xrightarrow[\left\|\cdot\right\|_{A}]{n\to\infty}\frac{\alpha p(A)g}{\left\|p(A)g\right\|_{A}}\,,$
where $\frac{\alpha p(A)g_{n}}{\left\|p(A)g\right\|_{A}}\in B_{V}$ for all $n\in\mathbb{N}$, $\alpha\in[-1,1]$. This shows us that $\mathcal{K}\left(A,g\right)\cap B_{V}\subset C$, and thus $\overline{\mathcal{K}\left(A,g\right)}^{V}\cap B_{V}\subset C$, which implies that $C=\overline{\mathcal{K}\left(A,g\right)}^{V}\cap B_{V}$ and therefore $\overline{\mathcal{K}\left(A,g_{n}\right)}^{V}\xrightarrow{\hat{d}^{V}_{w}}\overline{\mathcal{K}\left(A,g\right)}^{V}$. Moreover combining this with Proposition 3.11 gives us the convergence $\mathcal{K}\left(A,g_{n}\right)^{\perp_{V}}\xrightarrow{\hat{d}^{V}_{w}}\mathcal{K}\left(A,g\right)^{\perp_{V}}$.
We now prove statement (i). Due to the inclusions, we see that $A\mathcal{K}\left(A,g\right)^{\perp_{V}}\subset A\mathcal{K}\left(A,g_{n}\right)^{\perp_{V}}\subset\mathcal{K}\left(A,g_{n}\right)^{\perp}$ where the last inclusion is from Proposition 4.15. Owing to the convergence of the orthogonal complements (in $\mathcal{H}$) established in Corollary 5.5 and $\mathfrak{L}=\overline{\mathcal{K}\left(A,g\right)}$, taking the limit as $n\to\infty$ we see that
$A\mathcal{K}\left(A,g\right)^{\perp_{V}}\subset\mathcal{K}\left(A,g\right)^{\perp}\,,$
and therefore we have that $\mathscr{I}_{\mathcal{K}}\left(A,g\right)=\{0\}$. ∎
Remark 5.8.
We note that should we release the assumption of Theorem 5.7 that the Hamburger moment problem be determinate, we can still say something informative about the weak-gap convergence of the nested sequence of Krylov subspaces $(\overline{\mathcal{K}\left(A,g_{n}\right)}^{V})_{n\in\mathbb{N}}$ in $V$.
In such a case we have that $\overline{\mathcal{K}\left(A,g_{n}\right)}^{V}\subset\overline{\mathcal{K}\left(A,g_{n+1}\right)}^{V}\subset\mathfrak{L}\cap\mathcal{D}(A)\cong L^{2}(\mathbb{R},(1+\lambda^{2})\mu_{g}^{(A)})$ for all $n\in\mathbb{N}$. By Lemma 3.9 there exists $C\subset B_{V}$ weakly closed (in $V$) such that $\overline{\mathcal{K}\left(A,g_{n}\right)}\xrightarrow{\hat{d}^{V}_{w}}C$. It is also clear that $C\subset\mathfrak{L}\cap B_{V}=\{v\in\mathfrak{L}\cap\mathcal{D}(A)\,|\,\left\|v\right\|_{A}\leqslant 1\}$. Taking any $h\in L^{2}(\mathbb{R},(1+\lambda^{2})\mu_{g}^{(A)})$ we let $h_{n}(\lambda):=h(\lambda)\chi_{[-n,n]}(\lambda)$ so that $|h_{n}|\leqslant|h|$ and $h_{n}(A)g=h(A)g_{n}$ for all $n\in\mathbb{N}$. Then $\alpha\frac{\left\|h(A)g_{n}\right\|_{A}}{\left\|h(A)g\right\|_{A}}\leqslant 1$ for all $\alpha\in[-1,1]$ and due to Lebesgue Dominated Convergence we have
$\left\|h(A)g_{n}-h(A)g\right\|_{A}^{2}\xrightarrow{n\to\infty}0\,,$
so that clearly $\alpha\frac{h(A)g}{\left\|h(A)g\right\|_{A}}\in C$ for all $\alpha\in[-1,1]$. This gives the inclusion $\mathfrak{L}\cap B_{V}\subset C$. Therefore,
(5.3) $\overline{\mathcal{K}\left(A,g_{n}\right)}^{V}\xrightarrow{\hat{d}^{V}_{w}}\mathfrak{L}\cap\mathcal{D}(A)\,.$
6. Acknowledgements
The author wishes to acknowledge the support of the Italian National Institute for Higher Mathematics INdAM, and wishes to thank Marko Erceg (University of Zagreb, Croatia) and Alessandro Michelangeli (Prince Mohammad Bin Fahd University, Saudi Arabia, and University of Bonn, Germany) for many stimulating discussions.